### v2

# for Dynamic Neural Radiance Fields

Abstract
The accurate reconstruction of dynamic scenes with neural radiance fields is significantly dependent on the estimation of camera poses. Widely used structure-from-motion pipelines encounter difficulties in accurately tracking the camera trajectory when faced with separate dynamics of the scene content and the camera movement. To address this challenge, we propose DynaMoN. DynaMoN utilizes semantic segmentation and generic motion masks to handle dynamic content for initial camera pose estimation and statics-focused ray sampling for fast and accurate novel-view synthesis. Our novel iterative learning scheme switches between training the NeRF and updating the pose parameters for an improved reconstruction and trajectory estimation quality. The proposed pipeline shows significant acceleration of the training process. We extensively evaluate our approach on two real-world dynamic datasets, the TUM RGB-D and the BONN RGB-D Dynamic dataset. DynaMoN improves over the state-of-the-art both in terms of reconstruction quality and trajectory accuracy. We plan to make our code public to enhance research in this area.
I Introduction
Enabling novel view synthesis on dynamic scenes often requires multi-camera setups [1]. However, everyday dynamic scenes are often captured by one single moving camera, restricting the field of view [1]. In Simultaneous Localization and Mapping (SLAM) pipelines, multiple scene representations can be used. Commonly used explicit voxel and surfel representations are complemented by implicit truncated signed distance fields (TSDF) or neural radiance fields (NeRF). The latter allows the rendering of photorealistic novel views from new, reasonable camera positions [2, 3]. While many are restricted to static scenes, dynamic NeRF [4, 5, 6, 7] allows novel view synthesis for dynamic scenes.
To enable novel view synthesis using NeRF, accurate camera poses, usually retrieved via COLMAP [8] or the Apple ARKit, are essential. The structure-from-motion (SfM)-based approach often demands hours of computation to estimate reasonable camera poses and is challenged by large-scale dynamic scenes. To overcome the limitation of slow SfM approaches, recent works combined SLAM and NeRF [9, 10, 11, 2]. SLAM provides faster results for the camera trajectory than classic SfM. However, by nature, scenes are more often dynamic than static. Existing NeRF-SLAM [9, 10, 11, 2] methods assume a static scene which poses difficulties when dynamic scene content is present as shown in Fig 1.
Figure 1: NeRF-SLAM approaches rely on static scenes (left). Approaches such as InstantNGP [3] can be used to mask out dynamic scene content [2, 11] (left). Still, when using the masking, the scene representation lacks in quality (left). Fully dynamic NeRF provides an implicit 4D (3D+time) representation; however, it relies on offline SfM which can suffer in the presence of considerable motion (center). We consider dynamics via our motion-based segmentation (MS) and semantic segmentation (SS) already in the camera localization (right) and further during novel view synthesis. This enables a more robust camera tracking and novel view synthesis with a higher quality (ours, right).
In this paper, we present DynaMoN, a motion-aware camera localization and visualization approach which can handle highly dynamic scenes in camera pose retrieval and dynamic scene representation. We utilize semantic segmentation and generic motion masks during our camera localization process to enable robust tracking of the camera path in a dynamic environment. These masks are further utilized for ray sampling to train an efficient dynamic NeRF, based on HexPlane [4], that refines the initial camera poses, leading to improved reconstruction quality in scenes with strong motions.
We conduct experiments on two challenging datasets, namely the dynamic subset of the TUM RGB-D [12] and the BONN RGB-D Dynamic [13] dataset. The considered masking in DynaMoN allows us to outperform existing approaches in terms of the absolute trajectory error and novel view synthesis results on these datasets.
II Related Works
As DynaMoN is a combination of dynamic NeRF with fast and robust motion-aware camera localization, we group related work in topics for camera localization and their employment for novel view synthesis and neural representation for dynamic scenes.
II-A Camera Localization and Scene Dynamics
SfM and SLAM [14, 15, 16, 17] are arguably the two most common approaches for robust camera localization. More specifically, SLAM often targets real-time localization of a camera coupled with mapping. In addition to visual information, some approaches can utilize complementary sensory signals such as depth images and IMU data to further improve the accuracy. A common branch of work employs ORB descriptors to build a 3D map and track the camera [14, 17, 15].
In addition to classical or partially learnable SLAM approaches, Teed and Deng [16] introduce an end-to-end learnable method. Their approach, termed DROID-SLAM, consists of a dense bundle adjustment layer which enables recurrent iterative updates of camera poses and pixel-wise depth.
Typical camera localization approaches rely on a static scene assumption. However, natural scenes are often dynamic [18, 19, 20, 21]. To address this, Yu et al. [18] introduce a dynamic SLAM approach building upon ORB-SLAM2 [17]. Their DS-SLAM integrates a semantic segmentation network and a moving consistency check in the tracking process. Rünz et al. [22] enrich RGB-D SLAM with object-awareness and semantic segmentation capabilities while DynaSLAM [23] combines geometric and deep learning-based segmentation to mask dynamic objects in an attempt to ignore them for mapping. Others split the scene into dynamic and static parts. Dai et al. [20] build upon ORB-SLAM2 [17] and introduce a semantic segmentation using point correlation for this split and Ye et al. [21] divide the scene representation in a static flow field and a dynamic flow field.
Following this idea, we also split the scene into static parts and dynamic content using predicted segmentation masks.
II-B Camera Localization and NeRF
With the retrieved camera poses, a captured environment can be represented in 3D. Aside from explicit maps, an implicit scene representation can be established using a multilayer perceptron (MLP) [24]. Mildenhall et al. [24] introduce the use of an MLP as a neural representation for novel view synthesis, denoted as NeRF. The implicit representation is regressed by a photometric comparison of rendered views with input images. This allows the synthesis of photorealistic images from novel viewpoints at the cost of the requirement for accurate camera poses during training, typically captured by computationally intensive SfM with COLMAP [8]. An efficient alternative to SfM is the use of SLAM to retrieve camera poses.
In this regard, Sucar et al. [25] propose the usage of an MLP as scene representation in a SLAM framework and Zhu et al. [9] arrange multiple NeRFs in a hierarchical feature grid to learn both camera pose and scene representation. NICER-SLAM [10] introduces a similar approach for RGB input with estimated depth. Others purely rely on RGB and combine SLAM and NeRF [2, 11] utilizing hash-based InstantNGP [3] for scene representation. NeRF-SLAM [2] adapts the initial Droid-SLAM [16] framework and Orbeez-SLAM [11] leverages visual odometry [17].
To incorporate efficient pose estimation, we also rely on design choices of recent SLAM frameworks; however, we combine them with computationally efficient proposals [4] for neural radiance fields.
II-C Neural Representation of Dynamic Scenes
Our world is inherently dynamic which violates commonly used static scene assumptions [9, 10, 25, 11, 2, 3, 24]. To represent dynamics in a NeRF [26, 27], the use of grid structures became popular [4, 5]. HexPlane [4] utilizes a 4D space-time grid divided into six feature planes. The feature vector is a 4D point in space-time projected onto each feature plane. Similarly, Tensor4d [5] builds upon a 4D tensor field using time-aware volumes projected onto nine 2D planes. TiNeuVox [7] represents dynamic scenes using time-aware voxel features to enable faster training.
Optimizing for both camera poses and view synthesis can improve results in static scenes [28, 29, 30, 31, 32]. Liu et al. [33] therefore introduce such an optimization for a dynamic environment. To model this, they use one static and one dynamic NeRF.
We, in contrast, split the problem into differentiation of static and dynamic content already for camera pose estimation. The dynamic environment is then represented by a dynamic, 4D NeRF calculated upon the camera poses from the static scene components, which is further utilized to refine these poses.
III Method
We consider our input to be a video captured in a dynamic environment using a single monocular moving camera without pose knowledge. This leads to two objectives: first, estimating the camera pose, and second, reconstructing the scene which may comprise some deformation. We leverage masks to separate scene dynamics from static background for the former and utilize a 4D spatio-temporal neural representation for the latter. This allows to optimize for novel view synthesis while refining initial camera poses.
Figure 2: DynaMoN architecture. DynaMoN retrieves consecutive frames and applies semantic segmentation ($M_{ss}\rightarrow S_{S_{i}}$) and motion masks ($M_{MS}\rightarrow S_{M_{i}}$) on these input frames ($I$). Combining these masks (OR), we enable a motion-aware, fast and robust camera pose estimation. Based on the predicted camera poses $G$, we produce a time-dependent output ($\Theta$) using dynamic NeRF. During NeRF training, we utilize the previously computed masks to refine the camera poses (dashed lines). Dashed lines represent iterative updates on NeRF and pose parameters. We highlight our approach in green.
III-A Robust Camera Localization in Dynamic Environments
Novel view synthesis methods commonly employ rigid SfM, often COLMAP [8], for camera pose estimation. This assumes a static scene. To cope with dynamics, it is possible to add binary masks to remove non-static regions. However, the high computation time remains a drawback. In our approach, we target both robustness against dynamics and a more efficient solution for pose estimation. Our underlying architecture is based on DROID-SLAM [16]. While it incorporates elements to cope with a certain level of scene dynamics, we found it can be improved by eliminating the contribution of the non-static regions for camera tracking. We do this by generating binary dynamics masks for each frame, which are combinations of semantic ($M_{SS}$) and motions masks ($M_{MS}$).
The semantic mask follows the paradigm that in real videos, the most prominent source of motion is often an object of known class. We therefore define $M_{SS}=\mathbb{1}(x>0)\circ\sum_{i}M_{SS}(i)$ for all defined semantic classes $i\in S$ of the set $S$ with the indicator function being $1$ for positive values. Additional moving objects are covered through a motion mask $M_{MS}$. One example is depicted by the moving box in Fig 2. Following DytanVO [34], our motion masks are generated using a CNN with input of previous image, forward optical flow and estimated camera motion to iteratively refine the mask along with pose updates. Instead of random initialization, we begin with the motion mask of the previous frame to improve stability and lower the number of iterations in our refinement. We combine both masks as $M=\mathbb{1}(x>0)\left(M_{SS}+M_{MS}\right)$.
III-B Neural Dynamic Scenes and Pose Refinement
Our scene representation builds on one of the recent dynamic NeRF models, HexPlane [4]. Targeting efficient optimization, fast rendering and high reconstruction quality, HexPlane [4] utilizes TensoRF [35] and constructs a 4D representation with three spatial and one temporal axis to create feature planes with their pairwise combinations to be the input of their tiny MLP. The grid size of each axis is governed in a coarse-to-fine fashion along the training to improve convergence and naturally regularizes the local neighborhoods. For optimization, it follows the original NeRF’s ray casting formulation and applies a pixel-wise photometric loss between the rendered RGB values $C$ and the ground truth $\hat{C}$, supported by a total variation regularization as:
$\mathcal{L}=\frac{1}{|\mathcal{R}|}\sum_{\mathbf{r}\in\mathcal{R}}\|C(\mathbf{r})-\hat{C}(\mathbf{r})\|_{2}^{2}+\lambda_{\text{TV}}\mathcal{L}_{\text{TV}}.$ (1)
While $\mathcal{R}$ represents the sampled rays, $\mathcal{L}_{\text{TV}}$ and $\lambda_{\text{TV}}$ depict the total variation regularization and its weight.
TABLE I: RMSE of the translational ATE [m] on TUM RGB-D [12]. The upper rows are the slightly dynamic scenes and the lower rows are the highly dynamic scenes. The - denotes that no results were reported for these sequences.
Sequences RGB-D Monocular
DVO-SLAM  [36] ORB-SLAM2 [17] PointCorr [20] DynaPix-D [37] DytanVO [34] DeFlowSLAM [21] DROID-SLAM [16] OURS (MS&SS) OURS (MS)
fr2/desk-person 0.104 0.006 0.008 - 1.166 0.013 0.006 0.007 0.006
fr3/sitting-static 0.012 0.008 0.010 - 0.016 0.007 0.005 0.005 0.005
fr3/sitting-xyz 0.242 0.010 0.009 - 0.260 0.015 0.009 0.010 0.009
fr3/sitting-rpy 0.176 0.025 0.023 - 0.046 0.027 0.022 0.024 0.021
fr3/sitting-halfsphere 0.220 0.025 0.024 - 0.310 0.025 0.014 0.023 0.019
fr3/walking-static 0.752 0.408 0.011 0.007 0.021 0.007 0.012 0.007 0.014
fr3/walking-xyz 1.383 0.722 0.087 0.014 0.028 0.018 0.016 0.014 0.014
fr3/walking-rpy 1.292 0.805 0.161 0.123 0.155 0.057 0.040 0.031 0.039
fr3/walking-halfsphere 1.014 0.723 0.035 0.023 0.385 0.420 0.022 0.019 0.020
Mean/Max 0.577/1.383 0.304/0.805 0.041/0.161 - 0.265/1.166 0.065/0.420 0.016/0.040 0.016/0.031 0.016/0.039
III-C Pose Refinement and Training Strategy
We utilize eqn. (1) as the objective function for training both the NeRF and refining the camera poses. For the latter, we employ an alternating training strategy in which every 5,000 iterations the parameters to be optimized are switched between the radiance field and the pose parameters. To improve the accuracy and robustness of the optimization, during the pose refinement phase, we constrain the sampled rays, $\mathcal{R}$, to be elements of the static regions designated by the dynamics masks. Furthermore, we also use the same masks to ensure that the total variation regularization is imposed only on the static components. As the amount of non-dynamic pixels can vary from image to image, we accordingly scale the learning rate by the amount of static pixels in the image.
Figure 3: RMSE of the rotational ATE [°] on the TUM RGB-D dataset (left) and the BONN RGB-D Dynamic dataset (right). The median is shown in orange and the average in dotted lines.
IV Evaluation
DynaMoN is assessed on two challenging datasets, examining camera localization and novel view synthesis quality. To evaluate our approach, we consider two aspects. First, we consider the translational and rotational absolute trajectory error (ATE) [38] using root mean square error (RMSE) in meters and degrees. Second, we analyze the rendering quality of the novel dynamic views by reporting PSNR and SSIM [39] values.
TABLE II: RMSE of the translational ATE [m] on BONN RGB-D [13]. RGB-D results are reported by Palazzolo et al. [13]. For DROID-SLAM and ours, we calculated the results.
Sequences RGB-D Monocular
ReFusion [13] StaticFusion [40] DynaSLAM (G) [23] DynaSLAM (N+G) [23] DROID-SLAM [16] OURS (MS&SS) OURS (MS)
balloon 0.175 0.233 0.050 0.030 0.075 0.028 0.068
balloon2 0.254 0.293 0.142 0.029 0.041 0.027 0.038
balloon_tracking 0.302 0.221 0.156 0.049 0.035 0.034 0.036
balloon_tracking2 0.322 0.366 0.192 0.035 0.026 0.032 0.028
crowd 0.204 3.586 1.065 0.016 0.052 0.035 0.061
crowd2 0.155 0.215 1.217 0.031 0.065 0.028 0.056
crowd3 0.137 0.168 0.835 0.038 0.046 0.032 0.055
kidnapping_box 0.148 0.336 0.026 0.029 0.020 0.021 0.020
kidnapping_box2 0.161 0.263 0.033 0.035 0.017 0.017 0.017
moving_no_box 0.071 0.141 0.317 0.232 0.023 0.013 0.014
moving_no_box2 0.179 0.364 0.052 0.039 0.040 0.027 0.026
moving_o_box 0.343 0.331 0.544 0.044 0.177 0.152 0.167
moving_o_box2 0.528 0.309 0.589 0.263 0.236 0.175 0.176
person_tracking 0.289 0.484 0.714 0.061 0.043 0.148 0.024
person_tracking2 0.463 0.626 0.817 0.078 0.054 0.022 0.035
placing_no_box 0.106 0.125 0.645 0.575 0.078 0.021 0.018
placing_no_box2 0.141 0.177 0.027 0.021 0.030 0.020 0.020
placing_no_box3 0.174 0.256 0.327 0.058 0.025 0.022 0.022
placing_o_box 0.571 0.330 0.267 0.255 0.127 0.172 0.117
removing_no_box 0.041 0.136 0.016 0.016 0.016 0.015 0.015
removing_no_box2 0.111 0.129 0.022 0.021 0.020 0.019 0.021
removing_o_box 0.222 0.334 0.362 0.291 0.189 0.177 0.181
synchronous 0.441 0.446 0.977 0.015 0.006 0.007 0.006
synchronous2 0.022 0.027 0.887 0.009 0.012 0.006 0.011
Mean/Max 0.232/0.571 0.412/3.586 0.428/1.217 0.095/0.575 0.061/0.236 0.052/0.177 0.051/0.181
To assess DynaMoN, we compare both camera localization and novel view synthesis part with state-of-the-art approaches on two challenging datasets. For the novel view synthesis results, we follow the usual splitting using every 8th frame of the TUM RGB-D [12] dataset and BONN RGB-D Dynamic [13] dataset for testing.
IV-A Implementation Details
We implement our approach in Python using PyTorch. For semantic masking, we use DeepLabV3 [41] with a ResNet50-backbone [42] trained on the large-scale COCO dataset [43] and use human masks. For motion masking, we use the trained model of DytanVO [34]. In practice, initialized with the motion mask of the previous frame, we found that setting the number of iterations for motion mask refinement to two provides stable convergence. For cases where the camera motion is higher than in the training sequences of DytanVO, the motion segmentation module results in an increase of false positives; as a counter-measure, we introduce a threshold for the number of pixels that contain motion. Consequently, masks exceeding this threshold are discarded. This ensures that there are always enough pixels for the dense bundle adjustment.
While we use the reported image sizes for the novel view synthesis experiments, we use half of the resolution for the camera localization to be consistent with existent trajectory evaluations of DROID-SLAM [16].
Regarding the loss function, we set $\lambda_{TV}$ initially to 0.005 and the number of sampled rays $\lvert\mathcal{R}\rvert$ per batch to 1024.
To compare our camera pose retrieval with the classic way of using COLMAP, we follow the implementation of InstantNGP [3] to retrieve the COLMAP ground truth. Besides the COLMAP comparison, we compare our approach with RoDynRF [33]. For the TUM RGB-D dataset, we train their approach on an NVIDIA RTX 3090 with 24GB of VRAM. We follow their training guideline and do not employ prior pose initialization.
TABLE III: Novel view synthesis results on TUM RGB-D [12]. The upper rows are the slightly dynamic scenes and the lower rows are the highly dynamic scenes. The - denotes that the camera pose could not be regressed.
[…]
RoDynRF
HexPlane
HexPlane
+ DynaMoN
Backbone (Ours)
DynaMoN
fr2/desk-person 6.05 / 0.127 - 23.04 / 0.680 24.78 / 0.724
fr3/sitting-static 9.67 / 0.123 - 28.55 / 0.899 28.95 / 0.903
fr3/sitting-xyz 10.10 / 0.150 19.98 / 0.589 25.90 / 0.848 26.15 / 0.850
fr3/sitting-rpy 8.35 / 0.121 16.63 / 0.472 25.76 / 0.843 26.40 / 0.853
fr3/sitting-halfsphere 7.46 / 0.157 24.25 / 0.793 24.05 / 0.792 24.60 / 0.799
fr3/walking-static 10.01 / 0.159 20.27 / 0.650 25.84 / 0.849 26.25 / 0.853
fr3/walking-xyz 10.02 / 0.171 - 24.05 / 0.797 24.34 / 0.796
fr3/walking-rpy 8.81 / 0.201 - 24.18 / 0.790 24.70 / 0.796
fr3/walking-halfsphere 8.79 / 0.205 22.01 / 0.692 24.05 / 0.781 24.27 / 0.779
Mean 8.84 / 0.157 - 25.05 / 0.809 25.60 / 0.819
TABLE IV: Novel view synthesis results on BONN RGB-D Dynamic [13]. The - denotes that the camera pose could not be regressed.
[…]
HexPlane
HexPlane +
DynaMoN
Backbone (Ours)
DynaMoN
balloon - 29.83 / 0.868 29.84 / 0.866
balloon2 - 27.61 / 0.841 28.31 / 0.847
balloon_tracking - 29.27 / 0.847 29.55 / 0.846
balloon_tracking2 - 29.68 / 0.860 30.56 / 0.862
crowd - 27.92 / 0.831 28.40 / 0.829
crowd2 - 26.68 / 0.825 27.97 / 0.827
crowd3 - 26.57 / 0.815 27.40 / 0.827
kidnapping_box - 30.43 / 0.870 30.59 / 0.869
kidnapping_box2 30.87 / 0.881 30.28 / 0.860 30.18 / 0.857
moving_no_box 30.67 / 0.881 30.32 / 0.866 30.36 / 0.862
moving_no_box2 31.14 / 0.890 30.57 / 0.874 30.62 / 0.870
moving_o_box 29.21 / 0.889 30.26 / 0.884 31.08 / 0.888
moving_o_box2 - 29.82 / 0.875 29.40 / 0.875
person_tracking 26.57 / 0.808 27.78 / 0.843 28.71 / 0.845
person_tracking2 - 27.47 / 0.837 27.80 / 0.833
placing_no_box 25.73 / 0.780 30.94 / 0.869 31.23 / 0.870
placing_no_box2 - 31.18 / 0.885 31.45 / 0.884
placing_no_box3 - 30.79 / 0.879 30.83 / 0.875
placing_o_box - 30.42 / 0.886 31.12 / 0.890
removing_no_box - 30.66 / 0.879 30.85 / 0.877
removing_no_box2 31.13 / 0.887 30.73 / 0.871 30.86 / 0.870
removing_o_box 31.13 / 0.887 30.30 / 0.880 30.76 / 0.882
synchronous - 28.42 / 0.835 28.93 / 0.834
synchronous2 - 28.85 / 0.887 30.18 / 0.891
Mean - 29.45 / 0.861 29.87 / 0.861
Figure 4: Qualitative Results for Novel View Synthesis on TUM RGB-D dataset (top) and BONN RGB-D dynamic dataset (bottom). Shown are original Ground Truth images (left) and synthesized views from COLMAP-based dynamic NeRF (2nd column) with improved camera poses from our backbone (3rd column) only, and our full model (right).
IV-B Camera Localization Quality
Our results on all datasets show a lower translational trajectory error compared to state-of-the-art approaches, see Table I and Table II. We compare RGB-D-based and monocular SLAM approaches with our proposed camera retrieval method using only motion masks (MS) as well as using motion and semantic segmentation masks in combination (MS&SS). Looking at the mean value of the translational trajectory error, our approach performs similarly to DROID-SLAM on TUM RGB-D. However, the maximum error value is much lower, which shows that it is less prone to failure when applied to real-world sequences. Moreover, comparing the rotational and translational ATE, see Figure LABEL:fig:atel, our MS&SS achieves lower values than DROID-SLAM. On the BONN RGB-D Dynamic dataset, our approach outperforms the state-of-the-art approaches in terms of mean and maximum trajectory error.
TABLE V: Analysis of the influence of different pose retrieval steps to PSNR and SSIM in the Bonn RGB-D Dynamics crowd2 scene with sparse input trajectory.
Approach DROID-SLAM + HexPlane
Our initial poses + HexPlane
DynaMoN
PSNR $\uparrow$
[…]
16.46 / 0.661 17.07 / 0.660 17.31 / 0.659
IV-C Novel View Synthesis Quality
To evaluate our approach from a NeRF perspective we consider two aspects. First, the traditional way of generating camera ground truth when using NeRF, which is the use of COLMAP. Thus, we compare DynaMoN with HexPlane [4] using camera poses from COLMAP. Second, we highlight the influence of the iterative pose refinement by training both DynaMoN and HexPlane with our initially localized poses. Third, we compare our approach with RoDynRF [33] which regresses the camera pose along with a dynamic NeRF.
The results for the TUM RGB-D dataset are denoted in Table III. DynaMoN leads to a higher novel view synthesis quality compared to HexPlane without refined poses, RoDynRF, and when using COLMAP for the camera poses. Qualitative examples of our renderings are shown in Fig. 4.
On the BONN RGB-D dataset, COLMAP is challenged on several sequences (cf. Tab. IV). DynaMoN can retrieve camera poses for all scenes (Tab. II) and performs well on novel view synthesis for all sequences (Tab. IV). In addition, the PSNR is increased by the alternating training scheme. However, for the cases where COLMAP succeeds in the pose estimation, the camera poses from SfM also lead to decent novel view synthesis results. Examples of renderings can be seen in Fig. 4.
IV-D Robustness on Sparse Trajectories
Finally, we evaluate the robustness of different parts of our pose retrieval when faced with a dynamic and sparse trajectory. Using only every 10th frame of the crowd2 sequence, we show how the DROID-SLAM poses compare to the ones from the dynamics-masked camera pose retrieval and our neural camera pose refinement in terms of novel view rendering quality. Looking at the PSNR values in Table V, it can be noted that each of our proposed components incrementally increases the reconstruction quality. Thus, when confronted with a sparse trajectory of a highly dynamic scene, DynaMoN outperforms pure DROID-SLAM making it more robust to realistic, real-world scenarios with imperfect camera recordings.
V Discussion
In comparison to traditional sparse reconstructions, our approach addresses camera localization and novel view synthesis in combination. While this leads to more pleasing 3D visualizations and novel view synthesis results, the computing time is higher compared to traditional camera localization and 3D representation methods.
For camera pose estimation, our approach shows an improved performance compared to the state-of-the-art. Only using our motion masks already provides an improved performance. Moreover, for scenes with large motions, the combination with semantic segmentation masks provides significantly improved camera localization results. The results presented in Table III indicate that RoDynRF [33], despite being a state-of-the-art method that integrates combined camera regression with novel view synthesis, faces significant challenges in handling highly dynamic camera movements and dynamic scenes. In contrast, our method demonstrates superior robustness and yields improved outcomes when compared to both RoDynRF [33] and HexPlane [4] employing COLMAP poses on the demanding TUM RGB-D dataset. Furthermore, our iterative pose refinement process enhances robustness, as evidenced in Table V, and achieves better performance in novel view synthesis on both datasets.
VI Conclusion
We present DynaMoN, a motion-aware fast and robust camera localization approach for novel view synthesis. DynaMoN can handle not only the motion of known objects using semantic segmentation masks but also that of unknown objects using a motion segmentation mask. Furthermore, it retrieves camera poses faster and is more robust to scene dynamics in comparison to traditional SfM approaches, enabling a more accurate 4D scene representation. In addition, DynaMoN is trained with a novel iterative training scheme that refines NeRF parameters and the initially provided poses in an alternating fashion. Compared to the state-of-the-art, DynaMoN outperforms other dynamic camera localization approaches and shows better results for novel view synthesis on scenes with high camera motion.
Acknowledgment
We thank d.hip for providing a campus stipend.

### v3

@@
 Widely used structure-from-motion pipelines encounter difficulties in accurately tracking the camera trajectory when faced with separate dynamics of the scene content and the camera movement.
-To address this challenge, we propose DynaMoN.
+To address this challenge, we propose Dynamic Motion-Aware Fast and Robust Camera Localization for Dynamic Neural Radiance Fields (DynaMoN).
 DynaMoN utilizes semantic segmentation and generic motion masks to handle dynamic content for initial camera pose estimation and statics-focused ray sampling for fast and accurate novel-view synthesis.
@@
 The proposed pipeline shows significant acceleration of the training process.
-We extensively evaluate our approach on two real-world dynamic datasets, the TUM RGB-D and the BONN RGB-D Dynamic dataset.
+We extensively evaluate our approach on two real-world dynamic datasets, the TUM RGB-D dataset and the BONN RGB-D Dynamic dataset.
 DynaMoN improves over the state-of-the-art both in terms of reconstruction quality and trajectory accuracy.
@@
 Existing NeRF-SLAM [9, 10, 11, 2] methods assume a static scene which poses difficulties when dynamic scene content is present as shown in Fig 1.
+[^1] This work has been submitted to the IEEE for possible publication.
+Copyright may be transferred without notice, after which this version may no longer be accessible.
 Figure 1: NeRF-SLAM approaches rely on static scenes (left).
 Approaches such as InstantNGP [3] can be used to mask out dynamic scene content [2, 11] (left).
-Still, when using the masking, the scene representation lacks in quality (left).
-Fully dynamic NeRF provides an implicit 4D (3D+time) representation; however, it relies on offline SfM which can suffer in the presence of considerable motion (center).
+Still, when using the masking, the scene representation lacks quality (left).
+Fully dynamic NeRF provides an implicit 4D (3D+time) representation; however, it relies on offline SfM, which can suffer in the presence of considerable motion (center).
 We consider dynamics via our motion-based segmentation (MS) and semantic segmentation (SS) already in the camera localization (right) and further during novel view synthesis.
 This enables a more robust camera tracking and novel view synthesis with a higher quality (ours, right).
-In this paper, we present DynaMoN, a motion-aware camera localization and visualization approach which can handle highly dynamic scenes in camera pose retrieval and dynamic scene representation.
+In this paper, we present DynaMoN, a motion-aware camera localization and visualization approach that can handle highly dynamic scenes in camera pose retrieval and dynamic scene representation.
 We utilize semantic segmentation and generic motion masks during our camera localization process to enable robust tracking of the camera path in a dynamic environment.
@@
 The considered masking in DynaMoN allows us to outperform existing approaches in terms of the absolute trajectory error and novel view synthesis results on these datasets.
-II Related Works
+II RELATED WORK
 As DynaMoN is a combination of dynamic NeRF with fast and robust motion-aware camera localization, we group related work in topics for camera localization and their employment for novel view synthesis and neural representation for dynamic scenes.
@@
 In addition to classical or partially learnable SLAM approaches, Teed and Deng [16] introduce an end-to-end learnable method.
-Their approach, termed DROID-SLAM, consists of a dense bundle adjustment layer which enables recurrent iterative updates of camera poses and pixel-wise depth.
+Their approach, termed DROID-SLAM, consists of a dense bundle adjustment layer that enables recurrent iterative updates of camera poses and pixel-wise depth.
 Typical camera localization approaches rely on a static scene assumption.
@@
 Rünz et al.
-[22] enrich RGB-D SLAM with object-awareness and semantic segmentation capabilities while DynaSLAM [23] combines geometric and deep learning-based segmentation to mask dynamic objects in an attempt to ignore them for mapping.
+[22] enrich RGB-D SLAM with object-awareness and semantic segmentation capabilities, while DynaSLAM [23] combines geometric and deep learning-based segmentation to mask dynamic objects in an attempt to ignore them for mapping.
 Others split the scene into dynamic and static parts.
@@
 [21] divide the scene representation in a static flow field and a dynamic flow field.
+Yu et al.
+[24] combine semantic segmentation with optical flow to remove dynamic keypoints achieving a higher robustness compared to ORB-SLAM2.
 Following this idea, we also split the scene into static parts and dynamic content using predicted segmentation masks.
@@
 To model this, they use one static and one dynamic NeRF.
-We, in contrast, split the problem into differentiation of static and dynamic content already for camera pose estimation.
-The dynamic environment is then represented by a dynamic, 4D NeRF calculated upon the camera poses from the static scene components, which is further utilized to refine these poses.
+We, in contrast, split the problem into the differentiation of static and dynamic content already used for camera pose estimation.
+The dynamic environment is then represented by a dynamic, 4D NeRF calculated upon the camera poses from the static scene components, while further refining these poses based on the static parts of the images.
 III METHOD
-We consider our input to be a video captured in a dynamic environment using a single monocular moving camera without pose knowledge.
-This leads to two objectives: first, estimating the camera pose, and second, reconstructing the scene which may comprise some deformation.
-We leverage masks to separate scene dynamics from static background for the former and utilize a 4D spatio-temporal neural representation for the latter.
+We consider our input to be an RGB video captured in a dynamic environment using a single monocular moving camera without pose knowledge.
+This leads to two objectives: first, estimating the camera pose, and second, reconstructing the scene, which may comprise some deformation.
+We leverage masks to separate the scene dynamics from the static background for the former and utilize a 4D spatio-temporal neural representation for the latter.
 This allows to optimize for novel view synthesis while refining initial camera poses.
 Figure 2: DynaMoN architecture.
-DynaMoN retrieves consecutive frames and applies semantic segmentation ($M_{ss}\rightarrow S_{S_{i}}$) and motion masks ($M_{MS}\rightarrow S_{M_{i}}$) on these input frames ($I$).
+DynaMoN retrieves consecutive RGB frames and applies semantic segmentation ($M_{ss}\rightarrow S_{S_{i}}$) and motion masks ($M_{MS}\rightarrow S_{M_{i}}$) on these input frames ($I$).
 Combining these masks (OR), we enable a motion-aware, fast and robust camera pose estimation.
@@
 III-A Robust Camera Localization in Dynamic Environments
-Novel view synthesis methods commonly employ rigid SfM, often COLMAP [8], for camera pose estimation.
-This assumes a static scene.
+Novel view synthesis methods commonly employ rigid SfM, often COLMAP [8], for camera pose estimation. This assumes a static scene instead of dynamic scenes.
 To cope with dynamics, it is possible to add binary masks to remove non-static regions.
-However, the high computation time remains a drawback.
-In our approach, we target both robustness against dynamics and a more efficient solution for pose estimation.
+However, the high computation time of SfM approaches remains a drawback.
+In our approach, we target both robustness against dynamicscenes and a more efficient solution for pose estimation for dynamic moving cameras.
 Our underlying architecture is based on DROID-SLAM [16].
 While it incorporates elements to cope with a certain level of scene dynamics, we found it can be improved by eliminating the contribution of the non-static regions for camera tracking.
-We do this by generating binary dynamics masks for each frame, which are combinations of semantic ($M_{SS}$) and motions masks ($M_{MS}$).
-The semantic mask follows the paradigm that in real videos, the most prominent source of motion is often an object of known class.
-We therefore define $M_{SS}=\mathbb{1}(x>0)\circ\sum_{i}M_{SS}(i)$ for all defined semantic classes $i\in S$ of the set $S$ with the indicator function being $1$ for positive values.
-Additional moving objects are covered through a motion mask $M_{MS}$.
-One example is depicted by the moving box in Fig 2.
-Following DytanVO [34], our motion masks are generated using a CNN with input of previous image, forward optical flow and estimated camera motion to iteratively refine the mask along with pose updates.
-Instead of random initialization, we begin with the motion mask of the previous frame to improve stability and lower the number of iterations in our refinement.
-We combine both masks as $M=\mathbb{1}(x>0)\left(M_{SS}+M_{MS}\right)$.
-III-B Neural Dynamic Scenes and Pose Refinement
-Our scene representation builds on one of the recent dynamic NeRF models, HexPlane [4].
-Targeting efficient optimization, fast rendering and high reconstruction quality, HexPlane [4] utilizes TensoRF [35] and constructs a 4D representation with three spatial and one temporal axis to create feature planes with their pairwise combinations to be the input of their tiny MLP.
+We do this by generating binary dynamics masks for each frame, which are combinations of semantic ($M_{SS}$) and motion masks ($M_{MS}$).
+[…]
+III-B Neural Dynamic Scenes
+Our scene representation builds on a state-of-the-art dynamic NeRF, HexPlane [4].
+Targeting efficient optimization, fast rendering and high reconstruction quality, HexPlane [4] utilizes TensoRF [36] and constructs a 4D representation with three spatial and one temporal axes to create feature planes with their pairwise combinations to be the input of their tiny MLP.
 The grid size of each axis is governed in a coarse-to-fine fashion along the training to improve convergence and naturally regularizes the local neighborhoods.
-For optimization, it follows the original NeRF’s ray casting formulation and applies a pixel-wise photometric loss between the rendered RGB values $C$ and the ground truth $\hat{C}$, supported by a total variation regularization as:
+For optimization, we follow the original NeRF ray casting and apply the pixel-wise photometric loss between the rendered RGB values $C$ and the ground truth values $\hat{C}$, supported by a total variation regularization as:
 $\mathcal{L}=\frac{1}{|\mathcal{R}|}\sum_{\mathbf{r}\in\mathcal{R}}\|C(\mathbf{r})-\hat{C}(\mathbf{r})\|_{2}^{2}+\lambda_{\text{TV}}\mathcal{L}_{\text{TV}}.$ (1)
@@
 The - denotes that no results were reported for these sequences.
-Sequences RGB-D Monocular
-DVO-SLAM  [36] ORB-SLAM2 [17] PointCorr [20] DynaPix-D [37] DytanVO [34] DeFlowSLAM [21] DROID-SLAM [16] OURS (MS&SS) OURS (MS)
-fr2/desk-person 0.104 0.006 0.008 - 1.166 0.013 0.006 0.007 0.006
-fr3/sitting-static 0.012 0.008 0.010 - 0.016 0.007 0.005 0.005 0.005
-fr3/sitting-xyz 0.242 0.010 0.009 - 0.260 0.015 0.009 0.010 0.009
-fr3/sitting-rpy 0.176 0.025 0.023 - 0.046 0.027 0.022 0.024 0.021
-fr3/sitting-halfsphere 0.220 0.025 0.024 - 0.310 0.025 0.014 0.023 0.019
-fr3/walking-static 0.752 0.408 0.011 0.007 0.021 0.007 0.012 0.007 0.014
-fr3/walking-xyz 1.383 0.722 0.087 0.014 0.028 0.018 0.016 0.014 0.014
-fr3/walking-rpy 1.292 0.805 0.161 0.123 0.155 0.057 0.040 0.031 0.039
-fr3/walking-halfsphere 1.014 0.723 0.035 0.023 0.385 0.420 0.022 0.019 0.020
-Mean/Max 0.577/1.383 0.304/0.805 0.041/0.161 - 0.265/1.166 0.065/0.420 0.016/0.040 0.016/0.031 0.016/0.039
+Sequences RGB-D RGB
+DVO-SLAM ORB-SLAM2 PointCorr DynaPix-D [37] Yu et al. DytanVO DeFlowSLAM DROID-SLAM OURS (MS&SS) OURS (MS)
+[38] [17] [20] [37] [24] [35] [21] [16]
+fr2/desk-person 0.104 0.006 0.008 - - 1.166 0.013 0.006 0.007 0.006
+fr3/sitting-static 0.012 0.008 0.010 - 0.006 0.016 0.007 0.005 0.005 0.005
+fr3/sitting-xyz 0.242 0.010 0.009 - - 0.260 0.015 0.009 0.010 0.009
+fr3/sitting-rpy 0.176 0.025 0.023 - - 0.046 0.027 0.022 0.024 0.021
+fr3/sitting-halfsphere 0.220 0.025 0.024 - 0.018 0.310 0.025 0.014 0.023 0.019
+fr3/walking-static 0.752 0.408 0.011 0.007 0.009 0.021 0.007 0.012 0.007 0.014
+fr3/walking-xyz 1.383 0.722 0.087 0.014 0.014 0.028 0.018 0.016 0.014 0.014
+fr3/walking-rpy 1.292 0.805 0.161 0.123 0.031 0.155 0.057 0.040 0.031 0.039
+fr3/walking-halfsphere 1.014 0.723 0.035 0.023 0.020 0.385 0.420 0.022 0.019 0.020
+Mean/Max 0.577/1.383 0.304/0.805 0.041/0.161 - - 0.265/1.166 0.065/0.420 0.016/0.040 0.016/0.031 0.016/0.039
 III-C Pose Refinement and Training Strategy
@@
 (1) as the objective function for training both the NeRF and refining the camera poses.
-For the latter, we employ an alternating training strategy in which every 5,000 iterations the parameters to be optimized are switched between the radiance field and the pose parameters.
-To improve the accuracy and robustness of the optimization, during the pose refinement phase, we constrain the sampled rays, $\mathcal{R}$, to be elements of the static regions designated by the dynamics masks.
+For the latter, we employ an alternating training strategy in which every 5 000 iterations, the parameters to be optimized are switched between the radiance field and the pose parameters.
+To improve the accuracy and robustness of the optimization, during the pose refinement phase, we constrain the sampled rays, $\mathcal{R}$, contributing to the loss to be elements of the static regions designated by the dynamics masks.
 Furthermore, we also use the same masks to ensure that the total variation regularization is imposed only on the static components.
-As the amount of non-dynamic pixels can vary from image to image, we accordingly scale the learning rate by the amount of static pixels in the image.
+As the amount of non-dynamic sampled rays can vary from image to image, we scale the learning rate $\eta_{s}$ output by a learning rate scheduler by the fraction of static pixels in the image to keep the optimization steps in the same magnitude:
+$\eta=\eta_{s}*\frac{n_{static}}{n_{static}+n_{dynamic}},$ (2)
+where $n_{static}$ and $n_{dynamic}$ represent the number of static and dynamic pixels in the image and $\eta$ is the learning rate used for the pose refinement.
 Figure 3: RMSE of the rotational ATE [°] on the TUM RGB-D dataset (left) and the BONN RGB-D Dynamic dataset (right). The median is shown in orange and the average in dotted lines.
 IV EVALUATION
-DynaMoN is assessed on two challenging datasets, examining camera localization and novel view synthesis quality.
 To evaluate our approach, we consider two aspects.
@@
 Second, we analyze the rendering quality of the novel dynamic views by reporting PSNR and SSIM [40] values.
-TABLE II: RMSE of the translational ATE [m] on BONN RGB-D [13]. RGB-D results are reported by Palazzolo et al. [13]. For DROID-SLAM and ours, we calculated the results.
-Sequences RGB-D Monocular
+TABLE II: RMSE of the translational ATE [m] on BONN RGB-D Dynamic [13]. RGB-D results are reported by Palazzolo et al. [13]. For DROID-SLAM and ours, we calculated the results.
+Sequences RGB-D RGB
 ReFusion [13] StaticFusion [41] DynaSLAM (G) [23] DynaSLAM (N+G) [23] DROID-SLAM [16] OURS (MS&SS) OURS (MS)
@@
 Mean/Max 0.232/0.571 0.412/3.586 0.428/1.217 0.095/0.575 0.061/0.236 0.052/0.177 0.051/0.181
-To assess DynaMoN, we compare both camera localization and novel view synthesis part with state-of-the-art approaches on two challenging datasets.
-For the novel view synthesis results, we follow the usual splitting using every 8th frame of the TUM RGB-D [12] dataset and BONN RGB-D Dynamic [13] dataset for testing.
-IV-A Implementation Details
-We implement our approach in Python using PyTorch.
-For semantic masking, we use DeepLabV3 [41] with a ResNet50-backbone [42] trained on the large-scale COCO dataset [43] and use human masks.
-For motion masking, we use the trained model of DytanVO [34].
-In practice, initialized with the motion mask of the previous frame, we found that setting the number of iterations for motion mask refinement to two provides stable convergence.
-For cases where the camera motion is higher than in the training sequences of DytanVO, the motion segmentation module results in an increase of false positives; as a counter-measure, we introduce a threshold for the number of pixels that contain motion.
-Consequently, masks exceeding this threshold are discarded.
-This ensures that there are always enough pixels for the dense bundle adjustment.
-While we use the reported image sizes for the novel view synthesis experiments, we use half of the resolution for the camera localization to be consistent with existent trajectory evaluations of DROID-SLAM [16].
-Regarding the loss function, we set $\lambda_{TV}$ initially to 0.005 and the number of sampled rays $\lvert\mathcal{R}\rvert$ per batch to 1024.
-To compare our camera pose retrieval with the classic way of using COLMAP, we follow the implementation of InstantNGP [3] to retrieve the COLMAP ground truth.
-Besides the COLMAP comparison, we compare our approach with RoDynRF [33].
-For the TUM RGB-D dataset, we train their approach on an NVIDIA RTX 3090 with 24GB of VRAM.
-We follow their training guideline and do not employ prior pose initialization.
+IV-A Datasets
+To assess DynaMoN, we compare both the camera localization and the novel view synthesis part with state-of-the-art approaches on two challenging datasets.
+TUM RGB-D - Dynamic subset [12]
+The TUM RGB-D dataset has a resolution of $640\times 480$. While the dataset mainly focuses on static parts, five sequences were rated slightly dynamic and four highly dynamic.
+BONN RGB-D Dynamic [13]
+The BONN RGB-D Dynamic dataset consists of 24 dynamic sequences and two static sequences. The number of recordings per scene varies. The images have a resolution of $640\times 480$.
+For the novel view synthesis results, we follow the usual splitting using every eighth frame of the TUM RGB-D [12] dataset and BONN RGB-D Dynamic [13] dataset for testing.
+Figure 4: Qualitative Results for Novel View Synthesis on TUM RGB-D dataset (top) and BONN RGB-D Dynamic dataset (bottom).
+Shown are original Ground Truth images (left) and synthesized views from COLMAP-based dynamic NeRF (2nd column) with improved camera poses from our backbone (3rd column) only, and our full model (right).
 TABLE III: Novel view synthesis results on TUM RGB-D [12].
@@
 […]
-RoDynRF
+Sequences RoDynRF
 HexPlane
@@
 DynaMoN
-fr2/desk-person 6.05 / 0.127 - 23.04 / 0.680 24.78 / 0.724
+fr2/desk-person 6.05 / 0.127 - 23.04 / 0.680 24.83 / 0.725
 fr3/sitting-static 9.67 / 0.123 - 28.55 / 0.899 28.95 / 0.903
@@
 […]
+Sequences
 HexPlane
@@
 Mean - 29.45 / 0.861 29.87 / 0.861
-Figure 4: Qualitative Results for Novel View Synthesis on TUM RGB-D dataset (top) and BONN RGB-D dynamic dataset (bottom).
-Shown are original Ground Truth images (left) and synthesized views from COLMAP-based dynamic NeRF (2nd column) with improved camera poses from our backbone (3rd column) only, and our full model (right).
-IV-B Camera Localization Quality
+IV-B Implementation Details
+We implemented our approach in Python using PyTorch and evaluated it on one single workstation (Intel i9-10980XE CPU, NVIDIA GeForce RTX 3090 GPU 24GB, and 128GB RAM) unless otherwise specified.
+For semantic masking, we use DeepLabV3 [42] with a ResNet50-backbone [43] trained on the large-scale COCO dataset [44] and use human masks.
+For motion masking, we use the trained model of DytanVO [35].
+In practice, initialized with the motion mask of the previous frame, we found that setting the number of iterations for motion mask refinement to two provides stable convergence.
+For cases where the camera motion is higher than in the training sequences of DytanVO, the motion segmentation module results in an increase of false positives; as a counter-measure, we introduce a threshold of 60 % for the maximum amount of pixels that contain motion.
+Consequently, masks exceeding this threshold are discarded to ensure that there are always enough pixels for the dense bundle adjustment.
+Additionally, we only consider a pixel as dynamic if the confidence of the network exceeds 0.95 for the first and 0.98 for the remaining mask refinement steps.
+While we use the reported image sizes for the novel view synthesis experiments, we use half of the resolution for the camera localization to be consistent with existent trajectory evaluations of DROID-SLAM [16].
+Regarding the loss function, we set $\lambda_{TV}$ initially to 0.005 and the number of sampled rays $\lvert\mathcal{R}\rvert$ per batch to 1 024.
+We employ the default optimization settings for the dynamic NeRF from [4] and use a multi-step learning rate decay scheduler for the pose refinement.
+The initial learning rate for the scheduler is set to $1*10^{-5}$ at the beginning of each pose optimization cycle and decayed every 1 000 iterations by a factor of $0.9$.
+As described beforehand, the actual used learning rate of the optimizer is retrieved by multiplication with the percentage of static pixels in the image.
+The number of optimization iterations is set to 200 000, unless otherwise stated.
+To compare our camera pose retrieval with the classic way of using COLMAP, we follow the implementation of InstantNGP [3] to retrieve the COLMAP ground truth.
+Besides the COLMAP comparison, we compare our approach with RoDynRF [34].
+For the TUM RGB-D dataset, we train their approach on an NVIDIA RTX 3090 with 24GB of VRAM.
+We follow their training guideline and do not employ prior pose initialization.
+TABLE V: Analysis of the influence of different pose retrieval steps to PSNR and SSIM in the Bonn RGB-D Dynamic crowd2 scene with sparse input trajectory.
+Approach
+[…]
+DROID-SLAM [16] + HexPlane [4] 16.46 / 0.661
+Our initial poses + HexPlane [4] 17.07 / 0.660
+DynaMoN (Ours) 17.31 / 0.659
+TABLE VI: Analysis of the runtime on the TUM RGB-D fr2/desk-person scene. $\ddagger$ means 100 000 training iterations and a batch size of 512.
+$\dagger$ means 100 000 training iterations and a batch size of 1 024. The time format is hours:minutes:seconds.
+[…]
+RoDynRF [34] $\ddagger$ 01:52:55 08:00:44 00:11:15 10:04:54 6.05 / 0.127
+DynaMoN (Ours) 00:13:49 09:49:25 00:42:32 10:45:46 24.83 / 0.725
+DynaMoN (Ours) $\dagger$ 00:14:12 05:04:19 00:42:20 06:00:51 23.95 / 0.702
+DynaMoN (Ours) $\ddagger$ 00:13:41 04:43:23 01:03:21 06:00:25 23.43 / 0.690
+IV-C Camera Localization Quality
 Our results on all datasets show a lower translational trajectory error compared to state-of-the-art approaches, see Table I and Table II.
 We compare RGB-D-based and monocular SLAM approaches with our proposed camera retrieval method using only motion masks (MS) as well as using motion and semantic segmentation masks in combination (MS&SS).
-Looking at the mean value of the translational trajectory error, our approach performs similarly to DROID-SLAM on TUM RGB-D.
-However, the maximum error value is much lower, which shows that it is less prone to failure when applied to real-world sequences.
-Moreover, comparing the rotational and translational ATE, see Figure LABEL:fig:atel, our MS&SS achieves lower values than DROID-SLAM.
-On the BONN RGB-D Dynamic dataset, our approach outperforms the state-of-the-art approaches in terms of mean and maximum trajectory error.
-TABLE V: Analysis of the influence of different pose retrieval steps to PSNR and SSIM in the Bonn RGB-D Dynamics crowd2 scene with sparse input trajectory.
-Approach DROID-SLAM + HexPlane
-Our initial poses + HexPlane
-DynaMoN
-PSNR $\uparrow$
-[…]
-16.46 / 0.661 17.07 / 0.660 17.31 / 0.659
-IV-C Novel View Synthesis Quality
+Looking at the mean value of the RMSE of the translational trajectory error over all sequences, our approach performs similarly to DROID-SLAM on TUM RGB-D.
+However, the maximum error value is much lower, which shows that it is less prone to failure when applied to real-world scenarios.
+Moreover, comparing the RMSE of the rotational ATE, see Figure 3, our MS&SS-model achieves lower values than DROID-SLAM.
+On the BONN RGB-D Dynamic dataset, our approach outperforms the state-of-the-art approaches in terms of mean and maximum trajectory error for both rotation and translation.
+IV-D Novel View Synthesis Quality
 To evaluate our approach from a NeRF perspective we consider two aspects.
-First, the traditional way of generating camera ground truth when using NeRF, which is the use of COLMAP.
+First, the traditional way of generating camera ground truth when using NeRF is the use of COLMAP.
 Thus, we compare DynaMoN with HexPlane [4] using camera poses from COLMAP.
@@
 4.
-On the BONN RGB-D dataset, COLMAP is challenged on several sequences (cf.
+On the BONN RGB-D Dynamic dataset, COLMAP is challenged on several sequences (cf.
 Tab.
@@
 4.
-IV-D Robustness on Sparse Trajectories
-Finally, we evaluate the robustness of different parts of our pose retrieval when faced with a dynamic and sparse trajectory.
-Using only every 10th frame of the crowd2 sequence, we show how the DROID-SLAM poses compare to the ones from the dynamics-masked camera pose retrieval and our neural camera pose refinement in terms of novel view rendering quality.
+IV-E Robustness on Sparse Trajectories
+In this section, we evaluate the robustness of different parts of our pose retrieval when faced with a dynamic and sparse trajectory.
+Using only every tenth frame of the crowd2 sequence, we show how the DROID-SLAM poses compare to the ones from the dynamics-masked camera pose retrieval and our neural camera pose refinement in terms of novel view rendering quality.
 Looking at the PSNR values in Table V, it can be noted that each of our proposed components incrementally increases the reconstruction quality.
 Thus, when confronted with a sparse trajectory of a highly dynamic scene, DynaMoN outperforms pure DROID-SLAM making it more robust to realistic, real-world scenarios with imperfect camera recordings.
+IV-F Runtime and Memory Evaluation
+Finally, the runtime of DynaMoN is measured and the memory requirements are mentioned.
+Both are presented in comparison with RoDynRF [34] on a workstation with an NVIDIA RTX 4090 graphics card with 24 gigabytes of video random-access memory (VRAM).
+We report the results for the TUM RGB-D fr2/desk-person sequence in Table VI as it is by far the longest scene of both datasets comprising 4 067 frames.
+In terms of memory requirement, RoDynRF can only be trained with a batch size of 512 in this setting, while DynaMoN is more memory-efficient, enabling a batch size of 1 024.
+From the experiments, it can be concluded that the total time for pre-processing, training and rendering of DynaMoN is comparable to RoDynRF if the default training strategy with 200 000 iterations is utilized.
+However, if our model is trained for 100 000 iterations which is the default value of RoDynRF, the total time can be greatly reduced without a significant drop in photometric accuracy.
+For this shorter training, the NeRF and pose optimization of DynaMoN alternate every 2 500 iterations.
+Furthermore, a decrease of the batch size to the same value as in the competitor method further speeds up the training time at the cost of a longer rendering time while staying comparably accurate.
+It has to be noted that the pre-processing of RoDynRF consisting of monocular depth prediction, optical flow prediction, motion mask generation and dataloading takes about eight times longer than our pre-processing loading the data and generating initial poses as well as motion masks.
 V DISCUSSION
-In comparison to traditional sparse reconstructions, our approach addresses camera localization and novel view synthesis in combination.
-While this leads to more pleasing 3D visualizations and novel view synthesis results, the computing time is higher compared to traditional camera localization and 3D representation methods.
-For camera pose estimation, our approach shows an improved performance compared to the state-of-the-art.
-Only using our motion masks already provides an improved performance.
+In comparison to traditional sparse reconstructions [15, 14, 16], our approach addresses camera localization and dynamic novel view synthesis in combination.
+While this leads to more pleasing 3D visualizations and novel view synthesis results, the computing time is higher compared to traditional camera localization and non-dynamic 3D representation methods [9, 10, 11].
+Nevertheless, compared to combined dynamic camera localization and dynamic novel view synthesis approaches [34], our approach requires only 60% of the training time when using the same number of iterations and batch size, while being more robust in more dynamic scenes with dynamic camera motion.
+Moreover, our approach shows an improved performance compared to the state-of-the-art for camera localization.
+Only using our motion masks already improves the performance.
 Moreover, for scenes with large motions, the combination with semantic segmentation masks provides significantly improved camera localization results.

### v4

@@
-for Dynamic Neural Radiance Fields
+Alexander Ladikos4, Nassir Navab1,5, Daniel Roth2, and Benjamin Busam1
 Abstract
@@
 We plan to make our code public to enhance research in this area.
+Code available: https://hannahhaensen.github.io/DynaMoN/.
+Index Terms:
+Localization, Mapping
 I INTRODUCTION
-Enabling novel view synthesis on dynamic scenes often requires multi-camera setups [1].
+Enabling novel view synthesis (NVS) on dynamic scenes often requires multi-camera setups [1].
 However, everyday dynamic scenes are often captured by one single moving camera, restricting the field of view [1].
-In Simultaneous Localization and Mapping (SLAM) pipelines, multiple scene representations can be used.
-Commonly used explicit voxel and surfel representations are complemented by implicit truncated signed distance fields (TSDF) or neural radiance fields (NeRF).
+To receive NVS or a 3D/4D scene representation, accurate camera poses are essential for such a representation.
+One way to retrieve camera poses is Simultaneous Localization and Mapping (SLAM).
+SLAM can result in multiple scene representations such as explicit voxel or surfel representations or neural radiance fields (NeRF) representations.
 The latter allows the rendering of photorealistic novel views from new, reasonable camera positions [2, 3].
-While many are restricted to static scenes, dynamic NeRF [4, 5, 6, 7] allows novel view synthesis for dynamic scenes.
-To enable novel view synthesis using NeRF, accurate camera poses, usually retrieved via COLMAP [8] or the Apple ARKit, are essential.
-The structure-from-motion (SfM)-based approach often demands hours of computation to estimate reasonable camera poses and is challenged by large-scale dynamic scenes.
-To overcome the limitation of slow SfM approaches, recent works combined SLAM and NeRF [9, 10, 11, 2].
+While many are restricted to static scenes, dynamic NeRF [4, 5, 6, 7] allows NVS for dynamic scenes.
+To enable NVS, accurate camera poses, usually retrieved via structure-from-motion (SfM) [8] or the Apple ARKit, are essential.
+SfM often demands hours of computation to estimate reasonable camera poses and is challenged by large-scale dynamic scenes.
+To overcome the limitation of time-intensive SfM, recent works combined SLAM and NeRF [9, 10, 11, 2].
 SLAM provides faster results for the camera trajectory than classic SfM.
 However, by nature, scenes are more often dynamic than static.
-Existing NeRF-SLAM [9, 10, 11, 2] methods assume a static scene which poses difficulties when dynamic scene content is present as shown in Fig 1.
-[^1] This work has been submitted to the IEEE for possible publication.
-Copyright may be transferred without notice, after which this version may no longer be accessible.
-Figure 1: NeRF-SLAM approaches rely on static scenes (left).
-Approaches such as InstantNGP [3] can be used to mask out dynamic scene content [2, 11] (left).
-Still, when using the masking, the scene representation lacks quality (left).
+Existing Radiance Field-SLAM [9, 10, 11, 2, 12, 13] methods assume a static scene which poses difficulties when dynamic scene content is present as shown in Fig 1.
+Figure 1: NeRF combined with SLAM usually relies on static scenes (left).
+Approaches such as InstantNGP [3] can be used to mask out dynamic content [2, 11] (left) resulting still in reduced quality.
 Fully dynamic NeRF provides an implicit 4D (3D+time) representation; however, it relies on offline SfM, which can suffer in the presence of considerable motion (center).
-We consider dynamics via our motion-based segmentation (MS) and semantic segmentation (SS) already in the camera localization (right) and further during novel view synthesis.
-This enables a more robust camera tracking and novel view synthesis with a higher quality (ours, right).
-In this paper, we present DynaMoN, a motion-aware camera localization and visualization approach that can handle highly dynamic scenes in camera pose retrieval and dynamic scene representation.
-We utilize semantic segmentation and generic motion masks during our camera localization process to enable robust tracking of the camera path in a dynamic environment.
-These masks are further utilized for ray sampling to train an efficient dynamic NeRF, based on HexPlane [4], that refines the initial camera poses, leading to improved reconstruction quality in scenes with strong motions.
+DynaMoN considers dynamics at all stages using motion segmentation (MS) and semantic segmentation (SS) (right).
+This enables a more robust camera tracking and NVS with a higher quality (ours, right).
+In this paper, we present DynaMoN, a motion-aware camera localization and visualization approach that can handle highly dynamic scenes.
+  1. 1.
+    We propose a novel combination of semantic segmentation and generic motion masks during our camera localization to enable robust tracking of the camera path in a dynamic environment for dynamic neural radiance fields.
+  2. 2.
+    A statics-focused ray sampling building upon our dynamic masks is introduced to enhance our dynamic NeRF representation by refining the camera poses.
+Overall, (1) and (2) improve the camera poses and in turn improve the NVS, leading to improved NVS quality in scenes with strong motions.
 We conduct experiments on two challenging datasets, namely the dynamic subset of the TUM RGB-D [14] and the BONN RGB-D Dynamic [15] dataset.
-The considered masking in DynaMoN allows us to outperform existing approaches in terms of the absolute trajectory error and novel view synthesis results on these datasets.
+The considered masking in DynaMoN allows us to outperform existing approaches in terms of trajectory error and NVS results on these datasets.
 II RELATED WORK
-As DynaMoN is a combination of dynamic NeRF with fast and robust motion-aware camera localization, we group related work in topics for camera localization and their employment for novel view synthesis and neural representation for dynamic scenes.
+DynaMoN combines dynamic NeRF with fast and robust motion-aware camera localization.
+Therefore, related work covers camera localization, camera localization for NVS, and neural representation for dynamic scenes.
 II-A Camera Localization and Scene Dynamics
-SfM and SLAM [14, 15, 16, 17] are arguably the two most common approaches for robust camera localization.
-More specifically, SLAM often targets real-time localization of a camera coupled with mapping.
-In addition to visual information, some approaches can utilize complementary sensory signals such as depth images and IMU data to further improve the accuracy.
-A common branch of work employs ORB descriptors to build a 3D map and track the camera [14, 17, 15].
-In addition to classical or partially learnable SLAM approaches, Teed and Deng [16] introduce an end-to-end learnable method.
-Their approach, termed DROID-SLAM, consists of a dense bundle adjustment layer that enables recurrent iterative updates of camera poses and pixel-wise depth.
-Typical camera localization approaches rely on a static scene assumption.
-However, natural scenes are often dynamic [18, 19, 20, 21].
-To address this, Yu et al.
-[18] introduce a dynamic SLAM approach building upon ORB-SLAM2 [17].
-Their DS-SLAM integrates a semantic segmentation network and a moving consistency check in the tracking process.
-Rünz et al.
-[22] enrich RGB-D SLAM with object-awareness and semantic segmentation capabilities, while DynaSLAM [23] combines geometric and deep learning-based segmentation to mask dynamic objects in an attempt to ignore them for mapping.
-Others split the scene into dynamic and static parts.
-Dai et al.
-[20] build upon ORB-SLAM2 [17] and introduce a semantic segmentation using point correlation for this split and Ye et al.
-[21] divide the scene representation in a static flow field and a dynamic flow field.
-Yu et al.
-[24] combine semantic segmentation with optical flow to remove dynamic keypoints achieving a higher robustness compared to ORB-SLAM2.
-Following this idea, we also split the scene into static parts and dynamic content using predicted segmentation masks.
-II-B Camera Localization and NeRF
-With the retrieved camera poses, a captured environment can be represented in 3D.
-Aside from explicit maps, an implicit scene representation can be established using a multilayer perceptron (MLP) [25].
+Previous work on camera localization based on visual odometry (VO) [16], SfM [8] and SLAM [17, 18, 19, 20] extensively study architectures built for static scene assumptions.
+However, these methods show limited performance and often fail when they are encountered with the dynamic content commonly existent in real-world scenes.
+In recent years, various classical and learning-based methods have been introduced to induce and handle the dynamic motion of the scene for robust camera pose estimation.
+With an assumption of prior knowledge of the objects that can independently move in the scene, DS-SLAM [21] builds upon ORB-SLAM2 [20] and integrates a semantic segmentation network to mask out motion-prone content along with a moving consistency check in the tracking process.
+Similarly, Dai et al.
+[22] extend ORB-SLAM2 [20] to handle the dynamic content with sparse point-level dynamic motion segmentation estimated by the correlation of their individual transformations across the temporal dimension.
+Ye et al.
+[23] utilize dense optical flow and leverage a learning-based approach to estimate and mask out the content that moves independently of the rigid camera motion.
+Building a VO pipeline for highly dynamic stereo surgical videos, Hayoz et al.
+[24] combine background segmentation with an end-to-end learning-based method to estimate the contribution of each pixel’s motion for camera pose optimization that is structured in a differentiable training architecture.
+We build DynaMoN based on the findings of the prior work and utilize both the semantic segmentation [21] of motion-prone objects and dense learning-based motion segmentation [23] to mask out dynamic content within a SLAM architecture [19] to estimate camera pose.
+Furthermore, we leverage this information with our introduced static-focused ray sampling method to further fine-tune the camera poses during neural reconstruction.
+II-B Camera Localization and Radiance Fields
+Neural implicit methods gained popularity for scene representation.
 Mildenhall et al.
-[25] introduce the use of an MLP as a neural representation for novel view synthesis, denoted as NeRF.
-The implicit representation is regressed by a photometric comparison of rendered views with input images.
-This allows the synthesis of photorealistic images from novel viewpoints at the cost of the requirement for accurate camera poses during training, typically captured by computationally intensive SfM with COLMAP [8].
-An efficient alternative to SfM is the use of SLAM to retrieve camera poses.
-In this regard, Sucar et al.
-[26] propose the usage of an MLP as scene representation in a SLAM framework and Zhu et al.
-[9] arrange multiple NeRFs in a hierarchical feature grid to learn both camera pose and scene representation.
-NICER-SLAM [10] introduces a similar approach for RGB input with estimated depth.
-Others purely rely on RGB and combine SLAM and NeRF [2, 11] utilizing hash-based InstantNGP [3] for scene representation.
-NeRF-SLAM [2] adapts the initial Droid-SLAM [16] framework and Orbeez-SLAM [11] leverages visual odometry [17].
-To incorporate efficient pose estimation, we also rely on design choices of recent SLAM frameworks; however, we combine them with computationally efficient proposals [4] for neural radiance fields.
+[25] introduce NeRF employing a multi-layer perceptron (MLP) to represent a static 3D scene enabling NVS.
+NeRF requires precise camera poses during training, which are typically obtained through SfM pipelines like COLMAP [8] that can take long computation times for large sets of input images.
+Proposing an alternative solution that can jointly optimize camera poses and scene representation, Sucar et al.
+[26] reformulate NeRF in a SLAM framework for ordered sets of image sequences.
+Zhu et al.
+[9] extend this by arranging multiple NeRFs in a hierarchical structure, improving the reconstruction capabilities.
+Following this, NICER-SLAM [10] introduces the usage of estimated depth images for more accurate geometric reconstruction.
+Others [2, 11] utilize hash-based InstantNGP [3] for its fast novel view rendering capabilities.
+NeRF-SLAM [2] adapts the Droid-SLAM [19] framework and Orbeez-SLAM [11] leverages ORB-SLAM [20].
+More recent works [12, 27, 13] utilize an explicit scene representation for static scenes, namely Gaussian Splatting (GS) [28] combined with SLAM architectures for even faster convergence and novel view rendering.
+Unlike the prior work, our proposed method does not carry static-scene assumptions.
+To this end, we utilize a two-step approach that efficiently estimates the camera poses using a dynamics-aware SLAM method and represents the scene in a dynamic NeRF architecture [4] that is extended to further refine the initially estimated camera poses.
+Figure 2: DynaMoN retrieves consecutive RGB frames and applies semantic segmentation ($M_{ss}\rightarrow S_{S_{i}}$) and motion masks ($M_{MS}\rightarrow S_{M_{i}}$) on these input frames ($I$).
+Combining these masks (OR), we enable a motion-aware, fast and robust camera pose estimation $G$.
+Based on $G$, we produce a time-dependent output ($\Theta$) using dynamic NeRF.
+During NeRF training, we utilize the previously computed masks to refine the camera poses (dashed lines).
+Dashed lines represent iterative updates on NeRF and pose parameters.
+We highlight our approach in green.
 II-C Neural Representation of Dynamic Scenes
-Our world is inherently dynamic which violates commonly used static scene assumptions [9, 10, 26, 11, 2, 3, 25].
-To represent dynamics in a NeRF [27, 28], the use of grid structures became popular [4, 5].
+Our world is inherently dynamic which violates commonly used static scene assumptions and representations [9, 10, 26, 11, 2, 3, 29, 25].
+To represent dynamics [30, 31], the use of grid structures is common [4, 5, 7].
 HexPlane [4] utilizes a 4D space-time grid divided into six feature planes.
@@
 TiNeuVox [7] represents dynamic scenes using time-aware voxel features to enable faster training.
-Optimizing for both camera poses and view synthesis can improve results in static scenes [29, 30, 31, 32, 33].
-Liu et al.
-[34] therefore introduce such an optimization for a dynamic environment.
-To model this, they use one static and one dynamic NeRF.
-We, in contrast, split the problem into the differentiation of static and dynamic content already used for camera pose estimation.
-The dynamic environment is then represented by a dynamic, 4D NeRF calculated upon the camera poses from the static scene components, while further refining these poses based on the static parts of the images.
+[32, 33] propose the usage of GS with point cloud initializations for dynamic scene representations.
+Optimizing camera poses and view synthesis can improve results in static scenes [34, 35, 36, 29]. Liu et al. [37] introduce such an optimization for a dynamic NeRF using a static and a dynamic part.
+We, in contrast, split the problem into the differentiation of static and dynamic content.
+Already at the initial camera pose estimation this split is considered.
+The dynamic environment is then represented by an efficient dynamic, 4D NeRF [4] utilizing the camera poses estimated from the static scene components, and further refining these poses based on the static parts of the images.
 III METHOD
-We consider our input to be an RGB video captured in a dynamic environment using a single monocular moving camera without pose knowledge.
+We utilize robust and fast camera localization and an implicit scene representation, a dynamic NeRF, to enable NVS on dynamic scenes.
+As input, we consider an RGB video captured in a dynamic environment using a single monocular moving camera without pose knowledge.
 This leads to two objectives: first, estimating the camera pose, and second, reconstructing the scene, which may comprise some deformation.
 We leverage masks to separate the scene dynamics from the static background for the former and utilize a 4D spatio-temporal neural representation for the latter.
-This allows to optimize for novel view synthesis while refining initial camera poses.
-Figure 2: DynaMoN architecture.
-DynaMoN retrieves consecutive RGB frames and applies semantic segmentation ($M_{ss}\rightarrow S_{S_{i}}$) and motion masks ($M_{MS}\rightarrow S_{M_{i}}$) on these input frames ($I$).
-Combining these masks (OR), we enable a motion-aware, fast and robust camera pose estimation.
-Based on the predicted camera poses $G$, we produce a time-dependent output ($\Theta$) using dynamic NeRF.
-During NeRF training, we utilize the previously computed masks to refine the camera poses (dashed lines).
-Dashed lines represent iterative updates on NeRF and pose parameters.
-We highlight our approach in green.
+This allows to optimize for NVS while refining initial camera poses.
 III-A Robust Camera Localization in Dynamic Environments
-Novel view synthesis methods commonly employ rigid SfM, often COLMAP [8], for camera pose estimation. This assumes a static scene instead of dynamic scenes.
+NVS methods commonly employ rigid SfM, often COLMAP [8], for camera pose estimation which assumes a static scene.
 To cope with dynamics, it is possible to add binary masks to remove non-static regions.
@@
 […]
-III-B Neural Dynamic Scenes
-Our scene representation builds on a state-of-the-art dynamic NeRF, HexPlane [4].
-Targeting efficient optimization, fast rendering and high reconstruction quality, HexPlane [4] utilizes TensoRF [36] and constructs a 4D representation with three spatial and one temporal axes to create feature planes with their pairwise combinations to be the input of their tiny MLP.
+III-B Neural Dynamic Scene Representation
+To represent the dynamic scene, we build upon a state-of-the-art dynamic NeRF, HexPlane [4].
+Targeting efficient optimization, fast rendering and high reconstruction quality, HexPlane [4] utilizes TensoRF [39] and constructs a 4D representation with three spatial and one temporal axes to create feature planes.
+Their pairwise combinations are then the input of a tiny MLP.
 The grid size of each axis is governed in a coarse-to-fine fashion along the training to improve convergence and naturally regularizes the local neighborhoods.
@@
 While $\mathcal{R}$ represents the sampled rays, $\mathcal{L}_{\text{TV}}$ and $\lambda_{\text{TV}}$ depict the total variation regularization and its weight.
-TABLE I: RMSE of the translational ATE [m] on TUM RGB-D [12].
+III-C Pose Refinement and Training Strategy
+We utilize eqn.
+(1) as the objective function for training both the NeRF and refining the camera poses.
+During optimization, we employ an alternating training strategy in which every 5 000 iterations, the parameters to be optimized are switched between the radiance field and the pose parameters.
+To improve the accuracy and robustness of the optimization, during the pose refinement phase, we constrain the sampled rays, $\mathcal{R}$, contributing to the loss to be elements of the static regions designated by the inverse dynamics masks.
+Furthermore, we also use the same masks to ensure that the total variation regularization is imposed only on the static components.
+As the amount of non-dynamic sampled rays can vary from image to image, we scale the learning rate $\eta_{s}$ by the fraction of static pixels in the image to keep the optimization steps in the same magnitude:
+$\eta=\eta_{s}*\frac{n_{static}}{n_{static}+n_{dynamic}},$ (2)
+where $n_{static}$ and $n_{dynamic}$ represent the number of static and dynamic pixels in the image and $\eta$ is the learning rate used for the pose refinement.
+IV EVALUATION
+To evaluate our approach, we consider the translational absolute trajectory error (ATE) [40] using root mean square error (RMSE) in meters.
+To reason about the orientation, we employ the Relative Rotation Accuracy (RRA) as defined in [41] after synchronizing the poses to the GT to be invariant against the alignment of coordinate frames.
+We calculate the median RRA over all sequences and report the percentage of poses for which the RRA lies below a specified threshold.
+For NVS quality, we analyze the rendering quality of the novel dynamic views by reporting PSNR and SSIM [42].
+TABLE I: Percentage of orientations for which the median RRA is below the specified threshold [%] on the TUM RGB-D dataset and the Bonn RGB-D dynamic dataset.
+Dataset Metric DROID-SLAM OURS (MS&SS) OURS (SS) OURS (MS)
+TUM RGB-D median RRA@5 $\uparrow$ 100.0 100.0 100.0 100.0
+median RRA@15 $\uparrow$ 100.0 100.0 100.0 100.0
+Bonn RGB-D median RRA@5 $\uparrow$ 61.05 61.89 61.86 61.88
+median RRA@15 $\uparrow$ 91.29 91.45 91.46 91.27
+IV-A Datasets
+To assess DynaMoN, we evaluate both the camera localization and the NVS part with state-of-the-art approaches on two dynamic datasets.
+In total, we evaluate our approach on 33 dynamic sequences (9 from TUM-RGB-D, and 24 from BONN RGB-D).
+For NVS, we follow the usual splitting using every eighth frame for testing.
+TUM RGB-D - Dynamic subset [14]
+. We used the nine dynamic sequences, with five sequences rated slightly dynamic and four highly dynamic. The images of the TUM RGB-D dataset have a resolution of $640\times 480$.
+BONN RGB-D Dynamic [15]
+The BONN RGB-D Dynamic dataset consists of 24 dynamic sequences and two static sequences with an image resolution of $640\times 480$.
+TABLE II: RMSE of the translational ATE [m] on TUM RGB-D [14].
 The upper rows are the slightly dynamic scenes and the lower rows are the highly dynamic scenes.
@@
 Sequences RGB-D RGB
-DVO-SLAM ORB-SLAM2 PointCorr DynaPix-D [37] Yu et al. DytanVO DeFlowSLAM DROID-SLAM OURS (MS&SS) OURS (MS)
+DVO-SLAM ORB-SLAM2 PointCorr DynaPix-D [43] Yu et al. DytanVO DeFlowSLAM DROID-SLAM OURS (MS&SS) OURS (SS) OURS (MS)
 [44] [20] [22] [43] [45] [38] [23] [19]
-fr2/desk-person 0.104 0.006 0.008 - - 1.166 0.013 0.006 0.007 0.006
-fr3/sitting-static 0.012 0.008 0.010 - 0.006 0.016 0.007 0.005 0.005 0.005
-fr3/sitting-xyz 0.242 0.010 0.009 - - 0.260 0.015 0.009 0.010 0.009
-fr3/sitting-rpy 0.176 0.025 0.023 - - 0.046 0.027 0.022 0.024 0.021
-fr3/sitting-halfsphere 0.220 0.025 0.024 - 0.018 0.310 0.025 0.014 0.023 0.019
-fr3/walking-static 0.752 0.408 0.011 0.007 0.009 0.021 0.007 0.012 0.007 0.014
-fr3/walking-xyz 1.383 0.722 0.087 0.014 0.014 0.028 0.018 0.016 0.014 0.014
-fr3/walking-rpy 1.292 0.805 0.161 0.123 0.031 0.155 0.057 0.040 0.031 0.039
-fr3/walking-halfsphere 1.014 0.723 0.035 0.023 0.020 0.385 0.420 0.022 0.019 0.020
-Mean/Max 0.577/1.383 0.304/0.805 0.041/0.161 - - 0.265/1.166 0.065/0.420 0.016/0.040 0.016/0.031 0.016/0.039
-III-C Pose Refinement and Training Strategy
-We utilize eqn.
-(1) as the objective function for training both the NeRF and refining the camera poses.
-For the latter, we employ an alternating training strategy in which every 5 000 iterations, the parameters to be optimized are switched between the radiance field and the pose parameters.
-To improve the accuracy and robustness of the optimization, during the pose refinement phase, we constrain the sampled rays, $\mathcal{R}$, contributing to the loss to be elements of the static regions designated by the dynamics masks.
-Furthermore, we also use the same masks to ensure that the total variation regularization is imposed only on the static components.
-As the amount of non-dynamic sampled rays can vary from image to image, we scale the learning rate $\eta_{s}$ output by a learning rate scheduler by the fraction of static pixels in the image to keep the optimization steps in the same magnitude:
-$\eta=\eta_{s}*\frac{n_{static}}{n_{static}+n_{dynamic}},$ (2)
-where $n_{static}$ and $n_{dynamic}$ represent the number of static and dynamic pixels in the image and $\eta$ is the learning rate used for the pose refinement.
-Figure 3: RMSE of the rotational ATE [°] on the TUM RGB-D dataset (left) and the BONN RGB-D Dynamic dataset (right). The median is shown in orange and the average in dotted lines.
-IV EVALUATION
-To evaluate our approach, we consider two aspects.
-First, we consider the translational and rotational absolute trajectory error (ATE) [39] using root mean square error (RMSE) in meters and degrees.
-Second, we analyze the rendering quality of the novel dynamic views by reporting PSNR and SSIM [40] values.
-TABLE II: RMSE of the translational ATE [m] on BONN RGB-D Dynamic [13]. RGB-D results are reported by Palazzolo et al. [13]. For DROID-SLAM and ours, we calculated the results.
+fr2/desk-person 0.104 0.006 0.008 - - 1.166 0.013 0.006 0.007 0.007 0.006
+fr3/sitting-static 0.012 0.008 0.010 - 0.006 0.016 0.007 0.005 0.005 0.005 0.005
+fr3/sitting-xyz 0.242 0.010 0.009 - - 0.260 0.015 0.009 0.010 0.010 0.009
+fr3/sitting-rpy 0.176 0.025 0.023 - - 0.046 0.027 0.022 0.024 0.024 0.021
+fr3/sitting-halfsphere 0.220 0.025 0.024 - 0.018 0.310 0.025 0.014 0.023 0.028 0.019
+fr3/walking-static 0.752 0.408 0.011 0.007 0.009 0.021 0.007 0.012 0.007 0.007 0.014
+fr3/walking-xyz 1.383 0.722 0.087 0.014 0.014 0.028 0.018 0.016 0.014 0.014 0.014
+fr3/walking-rpy 1.292 0.805 0.161 0.123 0.031 0.155 0.057 0.040 0.031 0.036 0.039
+fr3/walking-halfsphere 1.014 0.723 0.035 0.023 0.020 0.385 0.420 0.022 0.019 0.019 0.020
+Mean/Max 0.577/1.383 0.304/0.805 0.041/0.161 - - 0.265/1.166 0.065/0.420 0.016/0.040 0.016/0.031 0.017/0.036 0.016/0.039
+TABLE III: RMSE of the translational ATE [m] on BONN RGB-D Dynamic [15]. RGB-D results are reported by Palazzolo et al. [15]. For DROID-SLAM and ours, we calculated the results.
 Sequences RGB-D RGB
-ReFusion [13] StaticFusion [41] DynaSLAM (G) [23] DynaSLAM (N+G) [23] DROID-SLAM [16] OURS (MS&SS) OURS (MS)
-balloon 0.175 0.233 0.050 0.030 0.075 0.028 0.068
-balloon2 0.254 0.293 0.142 0.029 0.041 0.027 0.038
-balloon_tracking 0.302 0.221 0.156 0.049 0.035 0.034 0.036
-balloon_tracking2 0.322 0.366 0.192 0.035 0.026 0.032 0.028
-crowd 0.204 3.586 1.065 0.016 0.052 0.035 0.061
-crowd2 0.155 0.215 1.217 0.031 0.065 0.028 0.056
-crowd3 0.137 0.168 0.835 0.038 0.046 0.032 0.055
-kidnapping_box 0.148 0.336 0.026 0.029 0.020 0.021 0.020
-kidnapping_box2 0.161 0.263 0.033 0.035 0.017 0.017 0.017
-moving_no_box 0.071 0.141 0.317 0.232 0.023 0.013 0.014
-moving_no_box2 0.179 0.364 0.052 0.039 0.040 0.027 0.026
-moving_o_box 0.343 0.331 0.544 0.044 0.177 0.152 0.167
-moving_o_box2 0.528 0.309 0.589 0.263 0.236 0.175 0.176
-person_tracking 0.289 0.484 0.714 0.061 0.043 0.148 0.024
-person_tracking2 0.463 0.626 0.817 0.078 0.054 0.022 0.035
-placing_no_box 0.106 0.125 0.645 0.575 0.078 0.021 0.018
-placing_no_box2 0.141 0.177 0.027 0.021 0.030 0.020 0.020
-placing_no_box3 0.174 0.256 0.327 0.058 0.025 0.022 0.022
-placing_o_box 0.571 0.330 0.267 0.255 0.127 0.172 0.117
-removing_no_box 0.041 0.136 0.016 0.016 0.016 0.015 0.015
-removing_no_box2 0.111 0.129 0.022 0.021 0.020 0.019 0.021
-removing_o_box 0.222 0.334 0.362 0.291 0.189 0.177 0.181
-synchronous 0.441 0.446 0.977 0.015 0.006 0.007 0.006
-synchronous2 0.022 0.027 0.887 0.009 0.012 0.006 0.011
-Mean/Max 0.232/0.571 0.412/3.586 0.428/1.217 0.095/0.575 0.061/0.236 0.052/0.177 0.051/0.181
-IV-A Datasets
-To assess DynaMoN, we compare both the camera localization and the novel view synthesis part with state-of-the-art approaches on two challenging datasets.
-TUM RGB-D - Dynamic subset [12]
-The TUM RGB-D dataset has a resolution of $640\times 480$. While the dataset mainly focuses on static parts, five sequences were rated slightly dynamic and four highly dynamic.
-BONN RGB-D Dynamic [13]
-The BONN RGB-D Dynamic dataset consists of 24 dynamic sequences and two static sequences. The number of recordings per scene varies. The images have a resolution of $640\times 480$.
-For the novel view synthesis results, we follow the usual splitting using every eighth frame of the TUM RGB-D [12] dataset and BONN RGB-D Dynamic [13] dataset for testing.
-Figure 4: Qualitative Results for Novel View Synthesis on TUM RGB-D dataset (top) and BONN RGB-D Dynamic dataset (bottom).
-Shown are original Ground Truth images (left) and synthesized views from COLMAP-based dynamic NeRF (2nd column) with improved camera poses from our backbone (3rd column) only, and our full model (right).
-TABLE III: Novel view synthesis results on TUM RGB-D [12].
+ReFusion [15] StaticFusion [46] DynaSLAM (G) [47] DynaSLAM (N+G) [47] DROID-SLAM [19] OURS (MS&SS) OURS (SS) OURS (MS)
+balloon 0.175 0.233 0.050 0.030 0.075 0.028 0.030 0.068
+balloon2 0.254 0.293 0.142 0.029 0.041 0.027 0.027 0.038
+balloon_tracking 0.302 0.221 0.156 0.049 0.035 0.034 0.037 0.036
+balloon_tracking2 0.322 0.366 0.192 0.035 0.026 0.032 0.029 0.028
+crowd 0.204 3.586 1.065 0.016 0.052 0.035 0.033 0.061
+crowd2 0.155 0.215 1.217 0.031 0.065 0.028 0.029 0.056
+crowd3 0.137 0.168 0.835 0.038 0.046 0.032 0.052 0.055
+kidnapping_box 0.148 0.336 0.026 0.029 0.020 0.021 0.020 0.020
+kidnapping_box2 0.161 0.263 0.033 0.035 0.017 0.017 0.017 0.017
+moving_no_box 0.071 0.141 0.317 0.232 0.023 0.013 0.014 0.014
+moving_no_box2 0.179 0.364 0.052 0.039 0.040 0.027 0.027 0.026
+moving_o_box 0.343 0.331 0.544 0.044 0.177 0.152 0.152 0.167
+moving_o_box2 0.528 0.309 0.589 0.263 0.236 0.175 0.176 0.176
+person_tracking 0.289 0.484 0.714 0.061 0.043 0.148 0.033 0.024
+person_tracking2 0.463 0.626 0.817 0.078 0.054 0.022 0.023 0.035
+placing_no_box 0.106 0.125 0.645 0.575 0.078 0.021 0.042 0.018
+placing_no_box2 0.141 0.177 0.027 0.021 0.030 0.020 0.019 0.020
+placing_no_box3 0.174 0.256 0.327 0.058 0.025 0.022 0.022 0.022
+placing_o_box 0.571 0.330 0.267 0.255 0.127 0.172 0.173 0.117
+removing_no_box 0.041 0.136 0.016 0.016 0.016 0.015 0.015 0.015
+removing_no_box2 0.111 0.129 0.022 0.021 0.020 0.019 0.020 0.021
+removing_o_box 0.222 0.334 0.362 0.291 0.189 0.177 0.179 0.181
+synchronous 0.441 0.446 0.977 0.015 0.006 0.007 0.007 0.006
+synchronous2 0.022 0.027 0.887 0.009 0.012 0.006 0.007 0.011
+Mean/Max 0.232/0.571 0.412/3.586 0.428/1.217 0.095/0.575 0.061/0.236 0.052/0.177 0.049/0.179 0.051/0.181
+IV-B Implementation Details
+We implemented our approach in Python using PyTorch and evaluated it on one single workstation (Intel i9-10980XE CPU, NVIDIA GeForce RTX 3090 GPU 24GB, and 128GB RAM) unless otherwise specified.
+For semantic masking, we use DeepLabV3 [48]/ResNet50 [49] trained on COCO [50] to obtain human masks.
+For motion masking, we apply DytanVO [38].
+In practice, initialized with the motion mask of the previous frame, we found that setting the number of iterations for motion mask refinement to two provides stable convergence.
+When the camera motion exceeds that of the DytanVO training sequences, the motion segmentation module produces more false positives.
+To counter this, we set a threshold of 60% for the maximum number of motion pixels; masks exceeding this threshold are discarded to maintain sufficient pixels for dense bundle adjustment.
+Additionally, a pixel is considered dynamic only if the confidence is above 0.95 for the first mask refinement step and 0.98 for subsequent steps.
+For NVS, we use the full image size. During camera localization, the images are downsampled by half to remain consistent with existing trajectory experiments [19].
+We set the loss parameter $\lambda_{TV}$ to 0.005 and sample $\lvert\mathcal{R}\rvert=1,024$ rays per batch.
+We use the default optimization settings for the dynamic NeRF [4], with a multi-step learning rate decay scheduler for pose refinement.
+The initial learning rate is $1\times 10^{-5}$, decayed by a factor of 0.9 every $1,000$ iterations and reset in each pose optimization cycle.
+The learning rate is scaled by the percentage of static pixels in the image.
+The number of optimization iterations is set to $200,000$, unless specified otherwise.
+IV-C Camera Localization Quality
+Our results on all datasets show a lower translational trajectory error compared to state-of-the-art approaches, see Table II and Table III.
+We compare RGB-D-based and monocular SLAM approaches with our proposed camera retrieval method using only motion masks (MS), using only semantic masks (SS) as well as using both types of masks in combination (MS&SS).
+Looking at the mean value of the RMSE of the translational ATE over all sequences, our approach performs similarly to DROID-SLAM on TUM RGB-D.
+However, the maximum error value is much lower, which shows that it is less prone to failure when applied to real-world scenarios.
+On the BONN RGB-D Dynamic dataset, our approach outperforms the state-of-the-art approaches in terms of mean and maximum trajectory error for translation.
+Moreover, comparing the RRA, see Table I, our MS&SS-model achieves higher values than DROID-SLAM on the BONN RGB-D dataset, while all methods perform equally well on TUM RGB-D.
+It can be noted that using one type of mask is usually sufficient to improve over DROID-SLAM.
+In some cases, like the person-tracking scene, where the combination performs worse than SS or MS, the SLAM backbone seems to have too less pixels left for the bundle adjustment.
+IV-D Novel View Synthesis Quality
+To evaluate the NVS quality of our approach, we consider several aspects.
+First, we compare it with the traditional way of generating camera ground truth for 4D NeRF [4] and 4D GS [32] using COLMAP, see Tables IV and V.
+We follow the implementation of InstantNGP [3] to retrieve the COLMAP ground truth.
+Second, we compare our approach with RoDynRF [37] without initial COLMAP poses, shown in Table IV.
+We follow RoDynRF’s training guidelines.
+Third, we highlight the influence of the iterative pose refinement by training both DynaMoN and HexPlane with our initially localized poses, see Tables  IV, Table V and  VI.
+Qualitative examples of renderings on both datasets can be found in Figure 3.
+On both datasets, COLMAP is challenged by the dynamics, and training with COLMAP camera poses can only be performed on a subset of scenes.
+DynaMoN retrieves all camera poses for both datasets, see Tables II and III.
+As visualized in Figure 3, the visual results of COLMAP-based NeRF are noisy on both datasets.
+DynaMoN’s camera poses can show improved results.
+The complete DynaMoN can further improve details.
+In terms of PSNR/SSIM, DynaMoN leads to a higher NVS quality compared to HexPlane without refined poses, RoDynRF, and when using COLMAP for the camera poses for either 4D NeRF or 4D GS, see Table IV.
+TABLE IV: NVS results on TUM RGB-D.
 The upper rows are the slightly dynamic scenes and the lower rows are the highly dynamic scenes.
 The - denotes that the camera pose could not be regressed.
+The * denotes random point cloud initialization, as the one from COLMAP lead to no convergence.
+We report PSNR/SSIM.
 […]
-Sequences RoDynRF
-HexPlane
+Sequences RoDynRF [37]
+4D Gaussians [32]
+HexPlane [4]
 HexPlane
@@
 DynaMoN
-fr2/desk-person 6.05 / 0.127 - 23.04 / 0.680 24.83 / 0.725
-fr3/sitting-static 9.67 / 0.123 - 28.55 / 0.899 28.95 / 0.903
-fr3/sitting-xyz 10.10 / 0.150 19.98 / 0.589 25.90 / 0.848 26.15 / 0.850
-fr3/sitting-rpy 8.35 / 0.121 16.63 / 0.472 25.76 / 0.843 26.40 / 0.853
-fr3/sitting-halfsphere 7.46 / 0.157 24.25 / 0.793 24.05 / 0.792 24.60 / 0.799
-fr3/walking-static 10.01 / 0.159 20.27 / 0.650 25.84 / 0.849 26.25 / 0.853
-fr3/walking-xyz 10.02 / 0.171 - 24.05 / 0.797 24.34 / 0.796
-fr3/walking-rpy 8.81 / 0.201 - 24.18 / 0.790 24.70 / 0.796
-fr3/walking-halfsphere 8.79 / 0.205 22.01 / 0.692 24.05 / 0.781 24.27 / 0.779
-Mean 8.84 / 0.157 - 25.05 / 0.809 25.60 / 0.819
-TABLE IV: Novel view synthesis results on BONN RGB-D Dynamic [13]. The - denotes that the camera pose could not be regressed.
+fr2/desk-person 6.05 / 0.127 - - 23.04 / 0.680 24.83 / 0.725
+fr3/sitting-static 9.67 / 0.123 - - 28.55 / 0.899 28.95 / 0.903
+fr3/sitting-xyz 10.10 / 0.150 26.40 / 0.886 19.98 / 0.589 25.90 / 0.848 26.15 / 0.850
+fr3/sitting-rpy 8.35 / 0.121 24.15 / 0.814 16.63 / 0.472 25.76 / 0.843 26.40 / 0.853
+fr3/sitting-halfsphere 7.46 / 0.157 24.42 / 0.830 24.25 / 0.793 24.05 / 0.792 24.60 / 0.799
+fr3/walking-static 10.01 / 0.159 16.39 / 0.575* 20.27 / 0.650 25.84 / 0.849 26.25 / 0.853
+fr3/walking-xyz 10.02 / 0.171 - - 24.05 / 0.797 24.34 / 0.796
+fr3/walking-rpy 8.81 / 0.201 - - 24.18 / 0.790 24.70 / 0.796
+fr3/walking-halfsphere 8.79 / 0.205 21.72 / 0.746 22.01 / 0.692 24.05 / 0.781 24.27 / 0.779
+Mean 8.84 / 0.157 - - 25.05 / 0.809 25.60 / 0.819
+TABLE V: NVS results on BONN RGB-D Dynamic. The - denotes that the camera pose could not be regressed. We report PSNR/SSIM.
 […]
@@
 Mean - 29.45 / 0.861 29.87 / 0.861
-IV-B Implementation Details
-We implemented our approach in Python using PyTorch and evaluated it on one single workstation (Intel i9-10980XE CPU, NVIDIA GeForce RTX 3090 GPU 24GB, and 128GB RAM) unless otherwise specified.
-For semantic masking, we use DeepLabV3 [42] with a ResNet50-backbone [43] trained on the large-scale COCO dataset [44] and use human masks.
-For motion masking, we use the trained model of DytanVO [35].
-In practice, initialized with the motion mask of the previous frame, we found that setting the number of iterations for motion mask refinement to two provides stable convergence.
-For cases where the camera motion is higher than in the training sequences of DytanVO, the motion segmentation module results in an increase of false positives; as a counter-measure, we introduce a threshold of 60 % for the maximum amount of pixels that contain motion.
-Consequently, masks exceeding this threshold are discarded to ensure that there are always enough pixels for the dense bundle adjustment.
-Additionally, we only consider a pixel as dynamic if the confidence of the network exceeds 0.95 for the first and 0.98 for the remaining mask refinement steps.
-While we use the reported image sizes for the novel view synthesis experiments, we use half of the resolution for the camera localization to be consistent with existent trajectory evaluations of DROID-SLAM [16].
-Regarding the loss function, we set $\lambda_{TV}$ initially to 0.005 and the number of sampled rays $\lvert\mathcal{R}\rvert$ per batch to 1 024.
-We employ the default optimization settings for the dynamic NeRF from [4] and use a multi-step learning rate decay scheduler for the pose refinement.
-The initial learning rate for the scheduler is set to $1*10^{-5}$ at the beginning of each pose optimization cycle and decayed every 1 000 iterations by a factor of $0.9$.
-As described beforehand, the actual used learning rate of the optimizer is retrieved by multiplication with the percentage of static pixels in the image.
-The number of optimization iterations is set to 200 000, unless otherwise stated.
-To compare our camera pose retrieval with the classic way of using COLMAP, we follow the implementation of InstantNGP [3] to retrieve the COLMAP ground truth.
-Besides the COLMAP comparison, we compare our approach with RoDynRF [34].
-For the TUM RGB-D dataset, we train their approach on an NVIDIA RTX 3090 with 24GB of VRAM.
-We follow their training guideline and do not employ prior pose initialization.
-TABLE V: Analysis of the influence of different pose retrieval steps to PSNR and SSIM in the Bonn RGB-D Dynamic crowd2 scene with sparse input trajectory.
+TABLE VI: Analysis of different pose retrieval steps on NVS quality, evaluated on the crowd2 scene of the BONN RGB-D Dynamic dataset.
+We assess robustness using a sparse input trajectory (every 10th frame).
 Approach
@@
 DynaMoN (Ours) 17.31 / 0.659
-TABLE VI: Analysis of the runtime on the TUM RGB-D fr2/desk-person scene. $\ddagger$ means 100 000 training iterations and a batch size of 512.
-$\dagger$ means 100 000 training iterations and a batch size of 1 024. The time format is hours:minutes:seconds.
-[…]
-RoDynRF [34] $\ddagger$ 01:52:55 08:00:44 00:11:15 10:04:54 6.05 / 0.127
-DynaMoN (Ours) 00:13:49 09:49:25 00:42:32 10:45:46 24.83 / 0.725
-DynaMoN (Ours) $\dagger$ 00:14:12 05:04:19 00:42:20 06:00:51 23.95 / 0.702
-DynaMoN (Ours) $\ddagger$ 00:13:41 04:43:23 01:03:21 06:00:25 23.43 / 0.690
-IV-C Camera Localization Quality
-Our results on all datasets show a lower translational trajectory error compared to state-of-the-art approaches, see Table I and Table II.
-We compare RGB-D-based and monocular SLAM approaches with our proposed camera retrieval method using only motion masks (MS) as well as using motion and semantic segmentation masks in combination (MS&SS).
-Looking at the mean value of the RMSE of the translational trajectory error over all sequences, our approach performs similarly to DROID-SLAM on TUM RGB-D.
-However, the maximum error value is much lower, which shows that it is less prone to failure when applied to real-world scenarios.
-Moreover, comparing the RMSE of the rotational ATE, see Figure 3, our MS&SS-model achieves lower values than DROID-SLAM.
-On the BONN RGB-D Dynamic dataset, our approach outperforms the state-of-the-art approaches in terms of mean and maximum trajectory error for both rotation and translation.
-IV-D Novel View Synthesis Quality
-To evaluate our approach from a NeRF perspective we consider two aspects.
-First, the traditional way of generating camera ground truth when using NeRF is the use of COLMAP.
-Thus, we compare DynaMoN with HexPlane [4] using camera poses from COLMAP.
-Second, we highlight the influence of the iterative pose refinement by training both DynaMoN and HexPlane with our initially localized poses.
-Third, we compare our approach with RoDynRF [34] which regresses the camera pose along with a dynamic NeRF.
-The results for the TUM RGB-D dataset are denoted in Table III.
-DynaMoN leads to a higher novel view synthesis quality compared to HexPlane without refined poses, RoDynRF, and when using COLMAP for the camera poses.
-Qualitative examples of our renderings are shown in Fig.
-4.
-On the BONN RGB-D Dynamic dataset, COLMAP is challenged on several sequences (cf.
-Tab.
-IV).
-DynaMoN can retrieve camera poses for all scenes (Tab.
-II) and performs well on novel view synthesis for all sequences (Tab.
-IV).
-In addition, the PSNR is increased by the alternating training scheme.
-However, for the cases where COLMAP succeeds in the pose estimation, the camera poses from SfM also lead to decent novel view synthesis results.
-Examples of renderings can be seen in Fig.
-4.
 IV-E Robustness on Sparse Trajectories
-In this section, we evaluate the robustness of different parts of our pose retrieval when faced with a dynamic and sparse trajectory.
-Using only every tenth frame of the crowd2 sequence, we show how the DROID-SLAM poses compare to the ones from the dynamics-masked camera pose retrieval and our neural camera pose refinement in terms of novel view rendering quality.
-Looking at the PSNR values in Table V, it can be noted that each of our proposed components incrementally increases the reconstruction quality.
+Dynamic scenes pose more challenges regarding the robustness of an approach compared to static environments.
+In Table VI, we showcase the robustness of our approach.
+We use only every tenth frame of the crowd2 sequence from the BONN RGB-D Dynamic dataset.
+As reported in Table VI, we show how the underlying camera localization approach influences the NVS result.
+When comparing DROID-SLAM camera poses to the ones from our DynaMoN backbone and our neural camera pose refinement, the best NVS result can be achieved using DynaMoN.
+Looking at the PSNR values in Table VI, it can be noted that each of our proposed components increases the reconstruction quality.
 Thus, when confronted with a sparse trajectory of a highly dynamic scene, DynaMoN outperforms pure DROID-SLAM making it more robust to realistic, real-world scenarios with imperfect camera recordings.
 IV-F Runtime and Memory Evaluation
-Finally, the runtime of DynaMoN is measured and the memory requirements are mentioned.
-Both are presented in comparison with RoDynRF [34] on a workstation with an NVIDIA RTX 4090 graphics card with 24 gigabytes of video random-access memory (VRAM).
-We report the results for the TUM RGB-D fr2/desk-person sequence in Table VI as it is by far the longest scene of both datasets comprising 4 067 frames.
+We compare the runtime and memory consumption of DynaMoN with RoDynRF [37] on a workstation with an NVIDIA RTX 4090 graphics card with 24 GB of VRAM.
+We report the results for the TUM RGB-D fr2/desk-person sequence in Table VII as it is by far the longest scene of both datasets comprising 4 067 frames.
+Using the COLMAP configuration from Instant-NGP, COLMAP runs for 38:28:16 (hours:minutes:seconds), see Table VII.
 In terms of memory requirement, RoDynRF can only be trained with a batch size of 512 in this setting, while DynaMoN is more memory-efficient, enabling a batch size of 1 024.
@@
 For this shorter training, the NeRF and pose optimization of DynaMoN alternate every 2 500 iterations.
-Furthermore, a decrease of the batch size to the same value as in the competitor method further speeds up the training time at the cost of a longer rendering time while staying comparably accurate.
+Furthermore, a decrease of the batch size to the same value as in the competitor method, RoDynRF, further speeds up the training time at the cost of a longer rendering time while staying comparably accurate.
 It has to be noted that the pre-processing of RoDynRF consisting of monocular depth prediction, optical flow prediction, motion mask generation and dataloading takes about eight times longer than our pre-processing loading the data and generating initial poses as well as motion masks.
+Figure 3: Qualitative Results for NVS on TUM RGB-D (top) and BONN RGB-D Dynamic (bottom).
+Ground Truth (left), novel views views from HexPlane + COLMAP (2nd column), improved camera poses from our backbone (3rd column), and our full model (right).
+TABLE VII: Analysis of the runtime on TUM RGB-D fr2/desk-person.
+Ours refers to 200,000 iterations / 1,024 batch size.
+The symbol $\ddagger$ indicates 100,000 iterations / 512 batch size, while $\dagger$ represents 100,000 iterations / 1.024 batch size.
+Time is given in the format hours:minutes:seconds
+[…]
+COLMAP [8] 38:28:16 - - - -
+RoDynRF [37] $\ddagger$ 01:52:55 08:00:44 00:11:15 10:04:54 6.05 / 0.127
+Ours 00:13:49 09:49:25 00:42:32 10:45:46 24.83 / 0.725
+Ours $\dagger$ 00:14:12 05:04:19 00:42:20 06:00:51 23.95 / 0.702
+Ours $\ddagger$ 00:13:41 04:43:23 01:03:21 06:00:25 23.43 / 0.690
 V DISCUSSION
-In comparison to traditional sparse reconstructions [15, 14, 16], our approach addresses camera localization and dynamic novel view synthesis in combination.
-While this leads to more pleasing 3D visualizations and novel view synthesis results, the computing time is higher compared to traditional camera localization and non-dynamic 3D representation methods [9, 10, 11].
-Nevertheless, compared to combined dynamic camera localization and dynamic novel view synthesis approaches [34], our approach requires only 60% of the training time when using the same number of iterations and batch size, while being more robust in more dynamic scenes with dynamic camera motion.
+In comparison to traditional sparse reconstructions [18, 17, 19], our approach addresses camera localization and dynamic NVS in combination.
+While this leads to more pleasing 3D visualizations and NVS results, the computing time is higher compared to traditional camera localization and non-dynamic 3D representation methods [9, 10, 11].
+Nevertheless, compared to combined dynamic camera localization and dynamic NVS approaches [37], our approach requires only 60% of the training time when using the same number of iterations and batch size, while being more robust in more dynamic scenes with dynamic camera motion.
 Moreover, our approach shows an improved performance compared to the state-of-the-art for camera localization.
@@
 Moreover, for scenes with large motions, the combination with semantic segmentation masks provides significantly improved camera localization results.
-The results presented in Table III indicate that RoDynRF [34], despite being a state-of-the-art method that integrates combined camera regression with novel view synthesis, faces significant challenges in handling highly dynamic camera movements and dynamic scenes.
-In contrast, our method demonstrates superior robustness and yields improved outcomes when compared to both RoDynRF [34] and HexPlane [4] employing COLMAP poses on the demanding TUM RGB-D dataset.
-Furthermore, our iterative pose refinement process enhances robustness, as evidenced in Table V, and achieves better performance in novel view synthesis on both datasets.
-VI CONCLUSION
-We present DynaMoN, a motion-aware fast and robust camera localization approach for novel view synthesis.
-DynaMoN can handle not only the motion of known objects using semantic segmentation masks but also that of unknown objects using a motion segmentation mask.
-Furthermore, it retrieves camera poses faster and is more robust to scene dynamics in comparison to traditional SfM approaches, enabling a more accurate 4D scene representation.
+The results presented in Table IV indicate that RoDynRF [37], despite being a state-of-the-art method that integrates combined camera regression with NVS, faces significant challenges in handling highly dynamic camera movements and dynamic scenes.
+In contrast, our method demonstrates superior robustness and yields improved outcomes when compared to both RoDynRF [37] and HexPlane [4] employing COLMAP poses on the BONN RGB-D Dynamic and TUM RGB-D dataset.
+Furthermore, our iterative pose refinement process enhances robustness, as evidenced in Table VI, and achieves better performance in NVS on both datasets.
+Looking at the results of 4D GS with COLMAP poses, combining robust camera localization and 4D GS seems to be a promising future research direction.
+VI LIMITATIONS
+DynaMoN is constrained by the scene representation capabilities of dynamic NeRFs.
+Although it performs well on large-scale datasets with up to over $4,000$ images, handling scenes with close to $8,000$ frames introduces significant challenges to the camera pose localization [51] and the NeRF [52].
+Nevertheless, our method works effectively with datasets like TUM RGB-D and BONN RGB-D Dynamic.
+Still, larger datasets [53] pose challenges, for NeRF and the camera localization [51].
+The underlying dynamic NeRF in this work has difficulties representing environments of this scale, limiting its applicability to more extensive datasets.
+VII CONCLUSION
+We present DynaMoN, a motion-aware, fast and robust camera localization approach for NVS.
+DynaMoN can cope with the motion of objects of known and unknown classes using the combined segmentation and motion mask resulting in the dynamics mask.
+DynaMoN retrieves camera poses faster and is more robust to scene dynamics in comparison to traditional SfM approaches, enabling a more accurate 4D scene representation.
 In addition, DynaMoN is trained with a novel iterative training scheme that refines NeRF parameters and the initially provided poses in an alternating fashion.
-Compared to the state-of-the-art, DynaMoN outperforms other dynamic camera localization approaches and shows better results for novel view synthesis on scenes with high camera motion.
-ACKNOWLEDGMENT
-We thank d.hip for providing a campus stipend.
+Compared to the state-of-the-art, DynaMoN outperforms other dynamic camera localization approaches and shows better results for NVS on scenes with high camera motion.